### v1

# Supplement to “Steering LLMs Towards Unbiased Responses: A Causality-Guided Debiasing Framework”

Abstract
Large language models (LLMs) can easily generate biased and discriminative responses. As LLMs tap into consequential decision-making (e.g., hiring and healthcare), it is of crucial importance to develop strategies to mitigate these biases. This paper focuses on social bias, tackling the association between demographic information and LLM outputs. We propose a causality-guided debiasing framework that utilizes causal understandings of (1) the data-generating process of the training corpus fed to LLMs, and (2) the internal reasoning process of LLM inference, to guide the design of prompts for debiasing LLM outputs through selection mechanisms. Our framework unifies existing de-biasing prompting approaches such as inhibitive instructions and in-context contrastive examples, and sheds light on new ways of debiasing by encouraging bias-free reasoning. Our strong empirical performance on real-world datasets demonstrates that our framework provides principled guidelines on debiasing LLM outputs even with only the black-box access.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs), which are trained on massive text corpora, have been found to exhibit concerning levels of social biases (Sheng et al., 2019; Gonen & Goldberg, 2019; Schick et al., 2021; Bender et al., 2021; Dodge et al., 2021). The unchecked biases can potentially perpetuate and amplify societal inequities, leading to unfair or even unethical outcomes. This issue is particularly significant as LLMs become more capable and start to serve as foundational components in decision-making systems across various sectors such as healthcare and education. Many debiasing approaches have been proposed to tackle this issue, for instance, direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023). For various reasons such as security and business interests, many of the most capable LLMs are closed-sourced, for instance, GPT-4 (Achiam et al., 2023), Gemini (Anil et al., 2023), Claude 2 (Anthropic, 2023), where the general public do not have access to models’ internal structures or parameters. Thus, prompting-based techniques largely become the only viable option to mitigate bias when using closed-sourced LLMs.
Figure 1: Different reasoning to answer a coreference question. A biased answer may be due to the use of a gender shortcut, while a bias-free answer is made by considering proper world knowledge given the circumstances. [^2] The use of demographic information does not necessarily indicate the reasoning is biased: sometimes certain demographic information (e.g., gender) should be considered for situations such as making medical decisions.
In this work, we focus on prompting techniques to steer LLMs towards unbiased responses. To study this, we notice that obtaining unbiased responses essentially boils down to the process of selecting proper pieces from the model’s internal representations and knowledge to make the decision. Let us consider a simple task of resolving the coreference of a gender pronoun in a given sentence. A model may output a biased answer by using a gender shortcut potentially learned from the imbalanced representations in the training data. For example, in Figure 2(a), the model may associate the pronoun “he” with “physician” rather than “secretary” due to social biases reflected in the training data – the gender distributions are disproportional for different occupations. We denote such reasoning process as biased reasoning, where the model selects internal representations in an improper way as this specific demographic information (gender) is unrelated to solving the coreference task. We provide more detailed discussions on selection mechanisms in Section 3.
Figure 2: Different prompting-based methods guided by one or both intuitions: discouraging ([-]) biased reasoning, and encouraging ([+]) bias-free reasoning.
Most of the existing prompting-based debiasing methods focus on discouraging biased reasoning, for instance, using explicit prompts to avoid using biased associations or prohibiting the utilization of gender information when making the decision. While these approaches help to a certain extent, another overlooked prompting-based strategy is to encourage bias-free reasoning by only selecting proper pieces related to the nature of the question. For example, Figure 2(b) demonstrates one possible way of conducting bias-free reasoning by comparing the likelihood of two situations occurring in real life and drawing on the more plausible situation to infer the coreference resolution.
We reveal the essential role of selection mechanisms in the interplay between LLM’s internal reasoning process and different strategies for designing the external prompts. We first construct a detailed causal model on the data generation process of the training corpus to identify ways biases may be smuggled in the pretraining phase. Then we construct a causal model of the LLM’s potential reasoning process and connect the two causal models by analyzing how different input prompts could modulate the LLM’s output through selection mechanisms. Building upon the above causal understandings, we introduce a causality-guided framework to debias LLMs, underpinned by the two intuitions: (1) discouraging biased reasoning  and (2) encouraging bias-free reasoning. Current prompting-based debiasing methods can also be viewed as instantiating one or both of the intuitions, as illustrated in Figure 2.
Using the causality-guided debiasing framework, we propose prompt-design strategies and further conduct a variety of empirical studies to analyze their effectiveness. We find that prompts employing both intuitions significantly outperform existing prompting-based debiasing methods. The strong empirical results clearly demonstrate the effectiveness of the causality-guided debiasing framework, even when we only have black-box accesses.
Our contributions can be summarized as follows:
  • We construct detailed causal modelings for both the underlying data-generating process of the training corpus and the LLM reasoning process. We note that selection mechanisms play an essential role in identifying how the LLM’s output could be modulated by different prompts.
  • We formulate a causality-guided debiasing framework. Through causal understandings of data-generating processes, we identify conditions that strategies for designing prompts should satisfy to debias in a principled and effective way. The framework also unifies existing prompting-based debias methods.
  • Using these strategies, we show strong empirical results on debiasing various social biases across different LLMs, demonstrating the clear benefit of our framework.
2 Preliminaries
In this section, we provide a brief introduction to causal modeling and causal reasoning (Section 2.1). We also provide a motivating example to illustrate how the selection mechanism can reshape dependence patterns within the different data (Section 2.2). [^3] We discuss related works in detail in Appendix A.
2.1 A Brief Introduction to Causality
For two random variables $X$ and $Y$, $X$ is a cause of $Y$ if there is a change in the distribution of $Y$ when we apply an intervention on $X$ while holding all other variables fixed (Spirtes et al., 1993; Pearl, 2009). We can represent causal relations among variables with a directed acyclic graph (DAG), where nodes represent variables, and edges represent direct causal relations between variables. We denote the direct causal relation between the ordered pair $(X,Y)$ by a directed edge $X\rightarrow Y$.
Local causal modules in a DAG, which characterize the causal relations between the corresponding variable and its direct causes, do not interfere with each other because of causal modularity. This property is also known as the exogeneity (Engle et al., 1983), or the independence of causal mechanism (Peters et al., 2017), resulting directly from the causal Markov condition for the DAG (Spirtes et al., 1993; Pearl, 2009). [^4] There are additional classes of graphs considered in causality literature, for example, directed cyclic graphs (DCGs) (Spirtes, 1995), ancestral graphs (Richardson & Spirtes, 2002), and so on. In this paper, we consider causal processes that can be modeled by a DAG. Other graph classes are beyond the scope of our work. In the context of language processing, the definition of a variable representing text or tokens is relatively abstract compared to the statistical notion of a random variable in tabular data. Within the scope of this work, we use the terms “variable” and “node” interchangeably when the context permits clear understandings.
2.2 A Motivating Example of Selection Mechanisms
Ideally, we would like samples to be drawn uniformly from the underlying population of interest. However, in practice, it is very common that the probability of including certain data points in the corpus depends on the characteristics of the data points themselves.
Previous literature has investigated selection mechanisms from different perspectives, for instance, the influence of selection bias on statistical inference in economic and sociological studies (Heckman, 1979, 1990; Winship & Mare, 1992), causal discovery when there are selection variables and latent common causes (Spirtes et al., 1995; Zhang, 2008), the identification and estimation of functional causal models when selection exists (Zhang et al., 2016), the identifiability of causal effect in the presence of selection bias from the graphical condition perspective (Bareinboim & Pearl, 2012; Bareinboim & Tian, 2015; Correa et al., 2019) and from the potential outcome perspective (Hernán & Robins, 2020), and the identification of the existence of selection bias from observational data under certain functional assumptions (Kaltenpoth & Vreeken, 2023).
Figure 3: An illustrative example of selection mechanisms.
Let us consider an example in the context of medical care. As illustrated in Figure 3, observed variables $(X_{1},X_{2})$ denote diseases, and $(Y_{1},Y_{2})$ denote corresponding symptoms. Apart from them, there exist potential binary selection variables $S_{i}$’s ($i\in\{1,2,\ldots,5\}$), where $S_{i}=1$ denotes being selected. $X_{1}$ ($X_{2}$) is the direct and only cause of $Y_{1}$ ($Y_{2}$), and the two diseases are unrelated in the general population, i.e., when none of the $S_{i}$’s exists. We use solid edges to represent causal relations among observed variables and dashed edges for those pertaining to selection mechanisms.
For instance, $S_{2}$ is an example of outcome-dependent selection (Zhang et al., 2016), which can be selecting individuals with symptom $Y_{1}$ from the general population. The conditional probability $P(Y_{1}\mid X_{1},S_{2}=1)$ in the selected data typically differs from its counterpart $P(Y_{1}\mid X_{1})$ in general population (Zhang et al., 2016). As another example, $S_{4}$ can denote the selection mechanism of only considering the hospital in-patient data. This signifies the setting of the well-known Berkson’s Paradox (Berkson, 1946), where two unrelated diseases $(X_{1},X_{2})$ appear to be correlated in the hospital data, simply because the data only contains selected patients who have at least one symptoms in $(Y_{1},Y_{2})$.
Figure 3 shows that the selection can be based solely on the cause (e.g., $S_{1}$), solely on the effect (e.g., $S_{2}$), or on both (e.g., $S_{3}$). Meanwhile, the selection can be based on variables that are not causally related in the general population (e.g., $S_{4}$ and $S_{5}$). Selection mechanisms can reshape dependence patterns among involved variables, and as a consequence, potentially change downstream outputs as well. We will see in Section 3 that such property of selection also applies to natural language processing (NLP) contexts.
(a) Training data corpus
(b) LLM reasoning process given prompt
(c) LLM reasoning process (annotated copy)
Figure 4: Causal graphs with respect to different data-generating processes. Panel (a) presents the underlying data-generating process of the training data corpus. Panels (b) – (c) present the reasoning process of the LLM. Since the conditions specified by the prompt are always conditioned upon when LLM generates the output, the internal reasoning is modulated by the external prompt. In panel (c) we include annotations to highlight causal pathways along which the information flow from demographic representation to LLM potential output is not regulated.
3 Causality-Guided Debiasing Framework
We propose our debiasing framework guided by causal understandings of involved data-generating processes. In Section 3.1, we present detailed causal models of both the underlying data-generating process for the training corpus and the LLMs’ reasoning process. We reveal that the LLM’s reasoning process is essentially an interplay between its internal representations and conditions specified by the external prompt designs, in which selection mechanisms play a key role. In Section 3.2, we identify different sets of conditions for prompt designs, which lead to prompting strategies that are not only intuitive but also theoretically grounded.
3.1 Detailed Causal Modeling of Processes
We present the causal modeling of the underlying data-generating process of the training data corpus (Section 3.1.1) and the reasoning process of LLMs (Section 3.1.2).
3.1.1 Underlying Data Generating Process of Text in Training Data Corpus
The training data corpus often reflects historical discrimination, and selection mechanisms are commonly involved. For instance, gender stereotype in occupations arises not due to the existence of a direct causal relation or a common cause between gender and occupation, but due to an underlying selection mechanism. Specifically, in the training data, among all possible combinations between gender (e.g., male and female) and occupation (e.g., CEO and secretary), there is a tendency to associate CEO more often with male, and secretary with female. This phenomenon occurs because the training data is a subset selected from an imaginary data corpus that is ideally diverse and comprehensive.
More generally, in Figure 4(a) we model causal relations in the underlying data-generating process of training data corpus (e.g., text scraped from the internet). Other than demographic information (e.g., race, gender, age), the causal graph contains additional variables of interest. We use “scenario” to represent the practical situation or context that sets the background of the text (e.g., medical care, hiring). We use “entity” to denote participants or stakeholders involved in the scenario, for instance, a patient may be an entity in the medical care scenario and a secretary may be an entity in the hiring scenario. The selection variable $S$ explicitly models the association between demographic information and entity, which is recognized as a major type of stereotype in NLP (Sweeney, 2013; Bolukbasi et al., 2016; Zhao et al., 2018; Tamkin et al., 2023). There are different types of texts in Figure 4(a). “Demographic-agnostic fact” denotes the text that does not explicitly contain demographic information. “Derivational text based on fact” denotes the text derived from fact, e.g., inference according to definition, fact-check Q&A, and restatement without altering factual contents. “Demographic-aware text” denotes the text where demographic information appears explicitly.
3.1.2 Reasoning Process of LLMs
Section 3.1.1 introduces causal modeling of the underlying data-generating process to characterize how discrimination is instantiated in LLMs’ training data corpus. The goal and expectation of the training is for LLM models to capture dependence patterns in the data corpus, and to extract and internalize the related knowledge. Therefore, we assume that the internal reasoning process of LLMs shares similarities with the underlying data-generating process of data corpus. Under this mild assumption, we formulate the reasoning process of how LLMs generate outputs based on input prompts in Figure 4(b). In particular, we reveal the interplay between internal representations and conditions specified by external inputs, and more importantly, how LLM outputs are shaped and modulated by different prompt designs through selection mechanisms.
In Figure 4(b), we use dotted contours to distinguish LLMs’ internal representations from the actual external information in the data corpus, e.g., the contrast between “demographic representation” in Figure 4(b) and “demographic information” in Figure 4(a). Notice that the internal nodes are not directly observable or accessible. We use double-stroke contours to indicate selection mechanisms, directed edges to represent direct causal relations, and dashed edges with hollow arrowheads to denote selection mechanisms.
A “prompt” serves as an external input to LLMs but not as a direct cause of internal representations. This is because the internal knowledge and representations exist beforehand, making them irrelevant to whether a specific prompt is provided. The prompt also does not act as an indicator for causal interventions. Because internal nodes are not directly observable or accessible, one cannot set them to certain values via hard interventions (Spirtes et al., 1993; Pearl, 2009; Peters et al., 2017; Hernán & Robins, 2020), or change the functional behavior of causal modules via soft interventions (Eberhardt & Scheines, 2007; Huang et al., 2020; Correa & Bareinboim, 2020). However, the prompt directly changes the selection variable “prompt properly considered” (PPC) through its designs. If the LLM model is well-trained and well-aligned, we can expect that PPC is always conditioned upon when the model produces the output, and the “LLM potential output” will become the actual output we obtain from the model. Just like the motivating example where selection can reshape dependence patterns among involved variables (Section 2.2), the prompt can influence the LLM reasoning process in significant ways by specifying conditions in selection mechanisms.
3.1.3 A Remark on the Modeled Processes
Figure 4(a) and Figure 4(b) each have their generating process of interest and distinct emphases. Although not pertaining to LLM reasoning itself, the underlying generating process modeled in Figure 4(a) provides hints on local causal modules of interest that prompt designs can specifically attend to for debiasing purposes. When we consider these hints with reference to the LLM reasoning process modeled in Figure 4(b), we can identify debiasing strategies that are both intuitive and theoretically grounded, as we will see in more detail in Section 3.2. The two detailed causal models complement each other, both of which are essential for understanding the source of bias, and furthermore, effectively debiasing LLMs.
3.2 Debiasing Guided by Causal Understandings
We present our prompt-based LLM debiasing framework guided by causal understandings of the related data-generating processes. Our core idea is to formulate conditions that should be specified in the prompt design, such that through the influence of selection mechanisms in the LLM reasoning process, one can effectively debias LLM outputs.
Figure 4(c) presents an annotated version of the LLM reasoning process (Section 3.1.2). Based on the understanding of the underlying generating process of training data corpus, and the mild assumption that the trained LLM captures the dependence patterns in training data, we highlight certain edges in light coral (accompanied by annotations marked with circled red numerals) in Figure 4(c) to denote unregulated information flow from demographic representations to LLM outputs in the internal reasoning process. We use circled blue Roman numerals to denote selection mechanisms that can be specified by the external input prompt (Section 3.1.2), and circled red Roman numerals to represent those corresponding to historical discrimination (Section 3.1.1).
According to the LLM reasoning process presented in Figure 4(c), we consider the information flows from the demographic representation (upstream node) to the LLM potential output (downstream node). We present additional conditions and constraints the prompt designs should specify for debiasing purposes. To clearly present the intuitions and theoretical groundings of our causality-guided framework, we identify three prompting strategies for debiasing LLMs, each of which serves as a solid starting point.
Strategy I (Nudge Towards Demographic-Agnostic Fact).
The intuition behind is to nudge LLMs towards utilizing demographic-agnostic fact when generating the output:
  1. Condition I
    The internal representations for demographic-agnostic fact and demographic information should be conditionally independent in the presence of PPC and existing selection $S$, i.e.,
[…]
Strategy I introduces Condition I to specify the selection mechanism over internal representations of demographic information and demographic-agnostic fact, denoted by edges i and v in Figure 4(c).
An example prompt employing Strategy I can be: “Considering the fact that the sentence ‘The physician hired the secretary because the secretary is highly recommended’ is practically more viable than the sentence ‘The physician hired the secretary because the physician is highly recommended’, who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
Strategy II (Counteract Existing Selection Bias).
The intuition behind this strategy is to directly counteract the effect of existing historical discriminations:
  1. Condition II.1
    The internal representations for demographic information and entity should be conditionally independent in the presence of PPC and existing selection $S$, i.e.,
[…]
  2. Condition II.2
    No new association between internal representations of demographic information and demographic-agnostic fact is introduced,
[…]
In Strategy II, Condition II.1 and Condition II.2 serve different purposes. Condition II.1 aims to counteract existing bias instantiated by selection $S$ (edges vi and vii) by constraining marginal dependence between representations for demographic information and entity (edges i and $i$).
Condition II.2 acts as a safeguard, making sure no new bias is introduced in the causal downstream of the entity so that the above counteraction effectively proceeds to the final output.
An example prompt employing Strategy II can be: “The physician can be either male or female, and the secretary can also be either male or female. Who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
Strategy III (Nudge Away from Demographic-Aware Text).
The intuition behind this strategy is to nudge LLMs away from utilizing demographic-aware text to generate outputs:
  1. Condition III
    The internal representations for demographic-aware text and demographic information should be conditionally independent in the presence of PPC and existing selection $S$,
[…]
Strategy III utilizes Condition III to specify the selection mechanism over internal representations of demographic information and demographic-aware text (edges i and iv) to regulate the information flow along the edge 1.
An example prompt employing Strategy III can be: “Do not use gender information to answer the question. Who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
Remarks on Three Strategies
The three strategies offer certain effectiveness individually but are not perfect on their own. [^5] We provide additional illustrations in Appendix B. While Strategy I nudges LLMs towards utilizing demographic-agnostic facts, it does not explicitly prevent LLMs from using demographic-aware text representations to generate the output, and the demographic information can potentially be associated to the output through an unregulated path containing edges 1 and 5. Similarly, while Strategy II aims to regulate information flows along edges { 2, 3, 4, vi, vii}, there is no explicit constraint involving edges 1 and 5, which leaves space for bias to sneak in during the reasoning process.
Compared to Strategy I that pushes LLMs to focus on demographic-agnostic facts, Strategy III prevents LLMs from referring to demographic-aware text during reasoning. While edge 1 is explicitly regulated by the selection mechanism specified by Condition III, the information flow along edge 5 can still result in the association between demographic information and the output. This is because Strategy III does not involve conditions or constraints over the dependence patterns among variables including demographic information and demographic-agnostic fact. As we can see from the conditions specified in Strategy II, no explicit appearance of demographic information in a demographic-agnostic fact does not guarantee the non-existence of dependence between the two. We would like to note that debiasing is better realized when the strategies are combined as they can address social bias in LLMs more comprehensively.
Table 1: Performance comparison of various debiasing methods on WinoBias. We show that combining encouraging bias-free reasoning and discouraging biased reasoning together (Reduce + Fact) significantly alleviates the gender bias for both coreference tasks: Type I coreference resolution requires world knowledge, and Type II coreference resolution can be resolved using only syntactic information. Pro in the table stands for coreference with pro-stereotypical pronouns, and Anti stands for coreference with anti-stereotypical pronouns. A smaller gap indicates less bias.
Accuracy (%) GPT 3 GPT 3.5 Claude 2 GPT 4
Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$
Type I
Default 43.01 79.24 36.23 62.96 94.03 31.07 67.57 92.13 24.56 82.50 97.96 15.47
COT (zero shot) 41.79 75.85 34.06 62.14 90.64 28.49 70.56 91.59 21.03 84.40 95.66 11.26
ICL 46.81 94.57 47.76 45.18 92.81 47.63 73.68 92.27 18.59 88.87 98.10 9.23
Reduce + Fact 73.27 73.95 0.68 72.73 84.67 11.94 74.08 75.17 1.09 94.57 96.74
Type II    Default 85.01 97.97 12.96 94.28 98.98 4.70 93.77 97.46 3.68 97.59 99.49 1.91 COT (zero shot) 69.12 86.79 17.66 93.90 98.35 4.45 95.30 99.62 4.32 97.97 99.62 1.65 ICL 94.41 99.62 5.21 95.43 98.98 3.56 94.03 97.84 3.81 98.35 99.87 1.52 Reduce + Fact 83.23 84.24 1.02 92.12 94.41 2.29 82.34 85.26 2.92 99.62 99.75 0.13
4 Results and Analyses
In this section, we demonstrate how the two types of approaches, i.e., (1) encouraging bias-free reasoning (Strategy I), and (2) discouraging biased reasoning (Strategy II and Strategy III), effectively steer LLMs towards unbiased response on various aspects of social bias (e.g., gender, race, and age). We present our experimental settings and results on the WinoBias data set (Zhao et al., 2018) in Section 4.1, and these on the Discrim-Eval data set (Tamkin et al., 2023) in Section 4.2.
4.1 Gender Bias: WinoBias
The WinoBias data set (Zhao et al., 2018) evaluates how likely models will assign stereotypical gender pronouns to occupations under coreference resolution tasks. The sentences in WinoBias are designed to be structurally parallel but differ in gender pronouns. It contains two sets of sentences: pro sentences with the pro-stereotypical gender pronouns (e.g., nurses as she, engineers as he), and anti sentences with anti-stereotypical gender pronouns (e.g., nurses as he, engineers as she). The data set also has two types of tasks with different levels of difficulties: coreference decisions in Type I task are challenging and must be made using world knowledge about given circumstances, whereas Type II task can be resolved using only syntactic information.
4.1.1 Experimental Settings (WinoBias)
For each sentence in WinoBias, we define the original question to be “Who does [gender pronoun] refer to in the sentence ‘[original sentence]’?”, and we measure the performance of four large language models: GPT-3, GPT-3.5, GPT-4, and Claude 2) across the above two types of coreference tasks (Type I and Type II).
Baseline Approaches
There are a few existing works employing prompting techniques to mitigate the bias in LLMs. We consider these three baselines:
  • Default: asks the original question directly without any additional prompts.
  • ICL with contrastive examples: provides contrastive examples followed by original question (Figure 2 middle). The examples are considered to be contrastive as the answer remains unchanged with different pronouns. In WinoBias, we use the same 16 ICL examples as in Si et al. (2022).
  • Zero-shot COT (Kojima et al., 2022): asks the model to also “think step by step” after the original question. Shaikh et al. (2022); Ganguli et al. (2023) design similar COT approaches to mitigate bias in LLMs.
Evaluation Metrics
We measure the performances of LLMs on pro and anti sentences in terms of accuracy, and the gap between the two indicates the level of gender bias exhibited by the models (a smaller gap is better).
Our Method: Reduce + Fact
We propose Reduce + Fact as one way of encouraging bias-free reasoning collectively with discouraging biased reasoning. As shown in Figure 2 right panel, we first create two gender-agnostic sentences by replacing the gender pronoun with the two occupations that appeared in the original sentence. Then we ask the model a factual question: which sentence is more likely to happen in real life, as this question does not contain any gender-related information. We then prepend its answer (to the factual question) before asking the original question. This allows us to distill LLM’s non-gender-related world knowledge (Strategy III) and nudge it to explicitly reason with this fact during coreference resolution (Strategy I). On top of encouraging bias-free reasoning (Strategies I and III), we also prompt that both occupations are equally likely to be male or female to counteract existing selection bias (Strategy II), discouraging biased reasoning.
One caveat of utilizing LLM’s world knowledge is that the performance of Reduce + Fact will increase as the capabilities of the LLM grow, since better world knowledge will further help LLM answer both factual and original questions.
4.1.2 Our Results (WinoBias)
We compare Reduce + Fact with the three baseline approaches on both coreference tasks and summarize our results in Table 3.2. We expect all LLMs to achieve higher accuracies on sentences containing pro-stereotypical pronouns (pro sentences) compared to these with anti-stereotypical pronouns (anti sentences) due to historical biases in the training data.
Coreference Resolution Requiring World Knowledge
For Type I tasks, world knowledge is required to perform coreference resolution. The Default prompting shows significant biases, with large gaps in accuracy between pro sentences and anti sentences for all models. The method Zero-shot COT marginally reduces the bias, as seen in the smaller gaps, but its performance is lower on both pro and anti scenarios for GPT-3 and GPT-3.5 when compared with Default, and it only marginally improves the general performance when applied to more capable ones (Claude 2 and GPT-4). For ICL with counterfactual examples, although the gap becomes larger for less capable models (GPT-3 and GPT-3.5), it can further reduce biases with more capable LLMs (Claude 2 and GPT-4), especially for GPT-4 (with a gap of 9.23%).
Remarkably, Reduce + Fact, which encourages bias-free reasoning while discouraging biased reasoning, substantially decreases the bias across all LLMs. The achieved minimal gaps, with GPT-4 exhibiting a mere 2.17% gap and 94.57% accuracy on anti sentences, demonstrate significantly greater effectiveness compared to other approaches.
Coreference Resolution Only with Syntactic Information
For Type II tasks, coreference can be resolved only with syntactic cues. All models generally show higher accuracies with smaller gaps. However, since there exists a shortcut in the Type II task – the correct answer is always the second entity in the sentence – a smaller gap may not necessarily indicate less bias. Still, Reduce + Fact outperforms the three baselines, particularly with GPT-4, achieving a near-negligible gap of 0.13%, suggesting that our method is highly effective in encouraging models to utilize non-gendered cues for coreference resolution.
Summary
These experimental results suggest prompt designs that encourage bias-free reasoning and discourage biased reasoning are effective at mitigating gender biases in LLMs by directing them to rely more on non-gender-related world knowledge (Strategies I and III) and less on gender shortcuts (Strategy II), thus promoting fairer and less biased responses. Also, the performance gap between pro sentences and anti sentences decreases as the LLMs become more capable, which may indicate that LLMs are less prone to assign occupations with stereotypical gender pronouns as their general (reasoning) capabilities grow.
Figure 5: Performance comparison on Discrim-Eval across three demographics. The bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Encouraging bias-free reasoning universally decreases the relative gap when added with methods that reduce biased reasoning.
4.1.3 Ablation studies
We further conduct ablation studies to individually examine the effectiveness of the two types of approaches, i.e., encouraging bias-free reasoning and discouraging biased reasoning. We divide the prompt Reduce + Fact into two parts: the Reduce Only part prompting both men and women can perform either occupation equally, and the Fact Only part telling the model which sentence(s) it regards to be more viable in real life.
Moreover, we categorize each LLM’s response into four groups: (i) TT: where the LLM answers both the factual question and the original question correctly, (ii) TF: where the LLM only answers the factual question correctly, (iii) FT: where the LLM only answers the original question correctly, and (iv) FF: where the LLM answers wrongly on both questions.
The above categorization allows us to attribute the model’s coreference mistakes into two categories: the mistake caused by its non-gender-related world knowledge (FF), and the mistakes caused by its gender bias (TF). Moreover, we can have a better understanding of the model’s correctness as well by looking at FT, where the correctness may be due to potentially biased shortcuts.
Table 2: Error Analysis on WinoBias Type I coreference task. We divide the models’ responses into 4 categories to better understand the cause of their errors and success. $F$ denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge. TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
GPT-4   Reduce + Fact 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Fact Only 91.45 95.93 5.29 0.14 0.54 0.81 2.71 3.12 Reduce Only 70.69 88.20 25.78 8.28 0.95 1.76 2.58 1.76 Default 69.47 87.25 26.87 8.68 1.36 2.44 2.31 1.63
As shown in Table 4.1.3, when we separate the prompting strategies into Fact Only and Reduce Only, we observe that Fact Only still performs quite well, but Reduce Only leads to a significant decrease in performance. This suggests that while discouraging biased reasoning is important, providing explicit instructions that promote bias-free reasoning is crucial for mitigating bias in LLMs.
Moreover, comparing the performance of GPT-4 with GPT-3.5, we can see that the GPT-4’s improvements mostly stem from better (non-gender-related) world knowledge (making way less mistakes on factual questions) and self-consistency (less mistakes in the TF group).
More experimental results can be found in Appendix D.1, where we design prompts to control the degree to which we steer LLMs away from biased reasoning.
4.2 Demographic Bias: Discrim-Eval
The Discrim-Eval data set (Tamkin et al., 2023) encompasses a collection of scenarios, each depicting a hypothetical case where a decision is required, such as approving a loan or issuing press credentials. The task is to make a Yes-or-No binary decision (affirmative or negative) on the individual involved in each scenario. The individuals can be characterized by three demographic factors: age (within a range of 20 to 100, in 10-year intervals), gender (male, female, or non-binary), and race (White, Black, Asian, Hispanic, or Native American).
4.2.1 Experimental Settings (Discrim-Eval)
The scenarios in the Discrim-Eval data set are designed such that an affirmative decision (Yes) is always favorable. The bias is measured using the probability of the Yes decision across a set of demographic variables (Tamkin et al., 2023).
Evaluation Metrics
For our evaluation of demographic bias, we use a relative gap to measure the adverse impact of different demographic variables (Hort & Sarro, 2022; Pessach & Shmueli, 2021). To be more specific, given a bias category that we want to evaluate (e.g., ethnicity), we obtain the probabilities of the ‘Yes’ token for all five different races, and we measure the gap between the largest probability and the smallest probability and divide the gap by the largest one. This way allows us to understand how substantial the discrimination is when comparing the most privileged group with the least privileged within the given bias category.
Baseline Approaches
The prompt-based approaches introduced in Tamkin et al. (2023) can all be viewed as ways to discourage biased reasoning, instantiating Strategy III. For our baseline, we choose the top four of them that reduce the discrimination the most (i.e., Illegal, Ignore, Really (4x), and Illegal + Ignore). The exact statements of each prompt can be found in Tamkin et al. (2023).
To amplify bias-free reasoning (Fact), we first remove all demographic information in each scenario and ask the LLM to make a Yes or No decision based on this base scenario. Then we explicitly include its answer to the base scenario in the prompt and ask the model to decide on the original scenario which contains specific demographic information (e.g., age 20).
4.2.2 Our Results (Discrim-Eval)
Since the evaluation metric requires accessing the logit of each token, our experiments on Discrim-Eval are conducted on these four LLMs: Mistral 7B, GPT-3, GPT-3.5, and GPT-4. Figure 5 shows our experimental results on three demographic categories: gender, ethnicity, and age.
Figure 6: GPT has an inherent bias towards aging. GPT-3’s default response (black) is more biased as age increases.
Each bar in the figure represents the relative gap, which measures the disparity in the probability of making an affirmative decision (‘Yes’) for the most and least privileged groups within each demographic category, characterizing the relative extent of discrimination. The bars are color-coded according to the methods used, where lighter shades correspond to the methods that explicitly encourage bias-free reasoning (e.g. purple bars denote Illegal and light purple bars denote Illegal + Fact). Default (black bar) denotes the prompt in the original scenario (with demographic information), without any modifications.
While different prompts of discouraging biased reasoning show various degrees of effectiveness in reducing demographic bias, no single prompt stands out as the universally most effective option across all LLMs and demographic categories. However, encouraging bias-free reasoning (Fact) collectively with discouraging biased reasoning can further decrease the relative gap, compared to employing the approaches on their alone. This observation is consistent across 4 LLMs and 3 demographic categories.
It is also noteworthy that none of the approaches achieves a zero gap. This indicates that while the debiasing prompt designs can reduce the discrimination in LLM outputs, they do not completely eliminate it.
Adverse Impact on Specific Demographic Groups
We further examine the relative gaps for specific demographic features. For example, in Figure 6, focusing on GPT-3 (which has the smallest relative gaps among the 4 chosen LLMs), we observe a growing trend of relative gaps on the default response (black bars) as the age increases. This suggests that GPT-3’s default responses may be more biased towards more senior individuals compared to junior ones. It is worth noting that the trend may potentially differ across LLMs, and we include additional plots for other LLMs on different demographic groups in Appendix D.2.
5 Conclusion
This paper presents a causality-guided and prompting-based LLM debiasing framework. In particular, we highlight the critical role of selection mechanisms in modeling data corpus bias and, more importantly, formulating how prompt designs can influence LLM outputs by specifying different selection conditions on its internal representations. Guided by causal understandings of such interplay, we identify principled prompting strategies for more effective debiasing. Our strong empirical results demonstrate the benefits of our framework, offering clear intuitions and theoretical foundations for effective debiasing approaches. Future work will naturally extend to acquiring bias-free knowledge and representations for LLMs.
Broader Impact
In this paper, we present a causality-based LLM output debiasing framework. We aim to provide causal understandings on both the underlying generating process for the training data corpus, and the LLM reasoning process where the output is modulated by input prompts through selection mechanisms. Our proposed framework serves as a general solution to promote fairness in pretrained LLMs without accesssing their model parameters. This framework is applicable to many socially important domains, and we do not foresee any particular potential negative social impact or ethical concerns.
Acknowledgement
This material is based upon work supported by the AI Research Institutes Program funded by the National Science Foundation (NSF) under AI Institute for Societal Decision Making (AI-SDM), Award No. 2229881. This project is also partially supported by an Amazon Research Award (Fall 2022 CFP), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Apple Inc., KDDI Research Inc., Quris AI, and Infinite Brain Technology.
Appendix A Related Works
In this section, we provide detailed discussions on related works. We consider the combinations of very related topics including causality, algorithmic fairness, and LLM reasoning. In Section A.1, we consider causal notions of fairness that do not specifically pertain to the LLM context. In Section A.2, we consider existing efforts to draw connections between causality and LLM reasoning. In Section A.3, we consider previous works on LLM debiasing. In Section A.4, we consider previous works that involve all three topics.
A.1 Causality and Fairness
It has been recognized in the algorithmic fairness literature that causality provides a unique tool to facilitate a better understanding of the data-generating process, and therefore, more effective bias quantifications and mitigations (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019; Nabi et al., 2022; von Kügelgen et al., 2022; Tang et al., 2023a). Previous causal fairness literature has considered notions based on estimating or bounding various kinds of causal effects (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019), and also the causal modeling of the dynamics as well as the long-term implications of bias mitigation strategies (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a, b).
Because of the relatively abstract definition of the variable or node in the language context, previous approaches for characterizing and enforcing causal fairness are not directly applicable in LLM debiasing tasks. That being said, as we have demonstrated in our causality-guided debiasing framework, causal understandings of the involved data-generating processes help identify effective debiasing strategies that are both intuitively clear and theoretically grounded.
A.2 Causality and LLMs
The intersection between causality and LLMs has drawn increasing attention. Zhang et al. (2023) considers three types of causal questions and aims to evaluate LLMs’ abilities to identify causal relations, discover new knowledge from data, and quantitatively estimate the consequences of actions. Kıcıman et al. (2023) investigate LLMs’ abilities to perform causal reasoning and solve covariance-/logic- based causal questions. They also study the failure modes of LLMs and provide techniques to interpret the model robustness. Jin et al. (2023) propose a benchmark data set for evaluating LLMs’ causal inference capabilities via the task of determining causal relationships from a set of correlational statements.
This line of research focuses on complex causal reasoning abilities in general settings, without specific attention to potential fairness violations. In comparison, our causality-guided debiasing framework does not involve assumptions/requirements on LLMs’ general-purpose causal reasoning capabilities. We adopt a rather mild assumption that a well-trained and well-aligned LLM captures the dependence pattern in the training data and that such a pattern is internalized and utilized during reasoning.
A.3 Debiasing Language Models
There is a large amount of work discussing bias and fairness in the context of language models (LMs) (Bordia & Bowman, 2019; Liang et al., 2021; Abid et al., 2021; Wang et al., 2023a; Liu et al., 2023; Ray, 2023; Rozado, 2023), and our investigation lies on debiasing techniques with causal understandings of the sources of biases. For debiasing approaches in the context of LMs, there are proposals involving direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), incorporating Reinforcement Learning with Human Feedback (RLHF) to better align the models with human values (Ouyang et al., 2022; Bai et al., 2022; Yao et al., 2023), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023). We focus on prompting-based techniques and identify principles for prompt designs to steer LLMs toward unbiased responses by a). reducing biased reasoning and b). encouraging bias-free reasoning. We provide demonstrations of how we can employ the above two principles. Works on LLMs reasoning such as Wei et al. (2022); Zheng et al. (2023) can also be incorporated into with our framework to encourage bias-free reasoning.
A.4 Debiasing LLMs from Causal Perspectives
Most closely related literature considers LLM debiasing strategies from causal perspectives. Vig et al. (2020) utilize causal mediation analysis and consider neuron-level intervention to investigate the instantiation of gender bias in Transformer-based language models (Vaswani et al., 2017). Zhou et al. (2023) propose Causal-Debias to mitigate the unwanted stereotypical association by fine-tuning pretrained language models. The causal model they consider involves four variables: label-relevant factor, bias-relevant factor, raw sentence, and ground-truth label. Wang et al. (2023b) pay special attention to entity bias and propose a specific structural causal model (SCM) for easier parameter estimations, such that the intervention-based mitigation strategy can be carried out. Their causal model involves four variables: entity, raw text, LLM input, and LLM output.
In comparison, we provide detailed causal modelings at a sub-sentence level, considering both the training corpus generating process and the LLM reasoning process. Furthermore, our framework explicitly models the interplay between internal representations and external inputs through selection mechanisms, providing a clearer picture regarding possible strategies to debias LLM outputs more effectively. Our approach does not require white-box access or the ability to perform interventions, making our prompting-based framework applicable to a variety of practical scenarios.
(a) LLM reasoning process (annotated with selection mechanisms specified by prompts)
(b) LLM reasoning process (Strategy I: nudge towards demographic-agnositic fact)
(c) LLM reasoning process (Strategy II: counteract existing selection bias)
(d) LLM reasoning process (Strategy III: nudge aware from demographic-aware text)
(e) LLM reasoning process (all strategies combined)
Figure 7: Additional illustrations on debiasing strategies.
Appendix B Further Illustrations and Discussions of Our Framework
In this section, we provide further illustrations on the debiasing strategies identified in Section 3.2.
In Figure 7, we use light coral (blue) to highlight unregulated (regulated) information flows from demographic representations to LLM outputs in the internal reasoning process. We use mix-colored highlight to denote partially regulated information flow, e.g., the edge from demographic-aware text representation to LLM potential output in Figure 6(d). Comparing Figure 6(e) (all strategies combined) with Figures 6(b)6(d) (strategies applied individually), we can see that collectively, the strategies address the social bias in language more comprehensively.
Here by “regulated”, we are referring to the constraint over the association between demographic information representation and the potential output. We would like to note that our causal models can naturally handle situations where certain kinds of neutral dependence patterns involving demographic information are not necessarily considered problematic (Tang et al., 2024). We present relevant experimental results in Section D.1. We would also like to note that our causality-guided framework is not limited to situations where demographic information is explicitly involved. One can adapt our framework to the specific practical setting by incorporating relevant nodes into causal graphs, thereby identifying the most suitable debiasing strategies therein.
Appendix C Experimental Details
Pretrained LLMs
For the GPT models used in our experiments, we consider snapshots from June 13th, 2023 where the knowledge cut-off time is Sep 2021. Since the legacy GPT-3 model (a.k.a., text-davinci-003) is no longer supported when we conduct the experiments, we use the model “gpt-3.5-turbo-instruct” instead as it has similar capabilities as GPT-3 era models. The Mistral-7B model we use in our experiments is the improved instruction fine-tuned version (a.k.a., “Mistral-7B-Instruct-v0.2”). We are unable to use Claude 2 for our experiments on the Discrim-Eval dataset because we cannot access the probabilities of generated tokens. All LLMs responses are obtained with a temperature of 0.
Iterative Prompting
The responses of LLMs may not follow a given format even with specific instructions, which makes answer extraction challenging. Moreover, we observe that adding instructions to restrict the format of responses may lower the quality of the generated outputs. Therefore, we apply 2-round iterative prompting in our experiments where we let the models generate freely and then ask them to summarize their answers in one or two words. This way allows us to obtain high answer qualities with an easy extraction process at the same time.
C.1 Gender Bias: WinoBias
For experiments on the WinoBias dataset, we combined both the training and test data for evaluation as there is no need to separate them when using prompting-based debiasing techniques. We also noticed that the original dataset contains 400 grammar mistakes across 3000+ sentences in total. We corrected these grammar mistakes as they may hinder the co-reference resolution tasks. We have provided the cleaned version in the supplementary materials.
Since there are 571 examples where two or more models make the same mistakes on the factual questions. After looking into these mistakes, we find that some of the examples are indeed ambiguous in nature (e.g., ”The construction worker stared at the teacher because [who] was angry”). Therefore, we collected human evaluations to understand the inherent ambiguity in the WinoBias dataset. For each of the 571 examples, we collect 5 human responses on the factual questions (i.e., which sentence is more likely to happen in real life), and we ask the human annotator to choose from {Sentence A, Sentence B, and Equally likely}. Based on the human responses, we identified 60 examples (55 from the Type I task and 5 from the Type II task) where three or more annotators disagree with the ground truth answers or think both sentences in the factual questions are equally likely to happen in real life. We removed these 60 examples during our evaluation, and we will include the human evaluation results in our codebase.
C.2 Demographic Bias: Discrim-Eval
The Discrim-Eval dataset contains 70 diverse decision scenarios and $9\times 3\times 5\times 70=9450$ individual decision questions which includes all combinations of [AGE] $\in$ [20, 30, 40, 50, 60, 70, 80, 90, 100], [GENDER] $\in$ [male, female, non-binary] and [RACE] $\in$ [white, Black, Asian, Hispanic, Native American]. To measure the corresponding bias in each demographic category, we reconstruct the dataset by extracting the base scenario which does not contain any demographic information (e.g., we replace all pronouns with the anaphoric reference to avoid leaking the gender information). We then ask the model to decide on each of the 70 base scenarios. There are (1/11/1/2) scenarios where (Mistral 7B/GPT-3/GPT-3.5/GPT-4) refuses to answer or does not output a Yes answer, and we removed these scenarios correspondingly when evaluating these LLMs.
Table 3: Error Analysis on WinoBias Type II coreference task. We divide the models’ responses into 4 categories to better understand their success and failure cases. $T$ denotes the (%) of examples where LLM answers both the factual question and the original question correctly, $F$ denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge. TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
GPT-4   Reduce + Fact 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25 Fact Only 99.36 99.87 0.25 0.00 0.13 0.00 0.25 0.13 Reduce Only 98.98 99.49 0.76 0.13 0.00 0.38 0.25 0.00 Default 98.73 99.75 0.89 0.13 0.25 0.00 0.13 0.13
Table 4: Error analysis of Reduce + Fact on Type I and Type II questions in WinoBias.
82.47 82.34 1.27 1.40 0.76 1.91 15.50 14.36   GPT-3.5   Type I 70.15 78.70 10.58 2.04 2.58 5.97 16.55 13.16 Type II 88.69 89.71 1.14 0.13 3.43 4.70 6.73 5.46   Claude   Type I 73.00 73.00 2.99 2.99 1.09 2.17 22.93 21.85 Type II 79.80 81.19 6.61 5.21 2.54 4.07 11.05 9.53   GPT-4   Type I 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Type II 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25
Table 5: Adjusting the levels of counteracting existing selection bias on GPT-3.5 with Reduce + Fact. We investigated how the level of counteracting stereotype impacts the model performance by adjusting the Fact part in the input prompt. For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time.
Appendix D Additional Experimental Results
D.1 Gender Bias: WinoBias
We include additional ablation studies on the WinoBias dataset, which include a detailed error analysis on WinoBias Type II coreference task (Table C.2), a detailed error analysis of Reduce + Fact with 4 LLMs on Type I and Type II tasks (Table C.2), and an ablation study on adjusting the levels of biased reasoning (Table C.2) in the prompt design.
In Section C.2, we show the detailed results on both Type I and Type II coreference tasks across 4 LLMs. As we can see, our method has bigger improvements on models with better world knowledge as. models with worse world knowledge could limit our method to reaching its full capacity.
Ablation study on adjusting the levels of counteracting existing selection bias
In Section C.2, we investigated how different levels of enforcing Strategy II (counteract existing selection bias) impacts the model debiasing performance by adjusting the Fact part in the input prompt. For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time; while a level of 50% indicates the following prompt: Assume that the physician can be male $50\%$ of the time and female $50\%$ of the time, and assume that the secretary can be male $50\%$ of the time and female $50\%$ of the time. We observe that as the level of anti-stereotype goes down, the errors caused by the use of the gender shortcut increase (TF increases). In addition, by soft adjustment of reducing biased reasoning, we provide not only flexible tuning strategies for the best model performance but also a chance to dive into the underlying reasons for the error.
D.2 Demographic Bias: Discrim-Eval
In Figure 8, Figure 9, Figure 10, Figure 11, we show in details the performance comparison on Discrim-Eval across three demographic categories across four LLMs (Mistral 7B, GPT 3, GPT 3.5, and GPT 4). The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). The baseline setting is colored black.
Figure 8: Performance comparison on Discrim-Eval across three demographic categories in details on Mistral (7B). The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
Figure 9: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-3. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
Figure 10: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-3.5. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
Figure 11: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-4. The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination). Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning). Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.

### v2

@@
-Supplement to “Steering LLMs Towards Unbiased Responses: A Causality-Guided Debiasing Framework”
+Prompting Fairness: Integrating Causality to Debias Large Language Models
 Abstract
-Large language models (LLMs) can easily generate biased and discriminative responses.
-As LLMs tap into consequential decision-making (e.g., hiring and healthcare), it is of crucial importance to develop strategies to mitigate these biases.
-This paper focuses on social bias, tackling the association between demographic information and LLM outputs.
-We propose a causality-guided debiasing framework that utilizes causal understandings of (1) the data-generating process of the training corpus fed to LLMs, and (2) the internal reasoning process of LLM inference, to guide the design of prompts for debiasing LLM outputs through selection mechanisms.
-Our framework unifies existing de-biasing prompting approaches such as inhibitive instructions and in-context contrastive examples, and sheds light on new ways of debiasing by encouraging bias-free reasoning.
-Our strong empirical performance on real-world datasets demonstrates that our framework provides principled guidelines on debiasing LLM outputs even with only the black-box access.
-Machine Learning, ICML
+Large language models (LLMs), despite their remarkable capabilities, are susceptible to generating biased and discriminatory responses.
+As LLMs increasingly influence high-stakes decision-making (e.g., hiring and healthcare), mitigating these biases becomes critical.
+In this work, we propose a causality-guided debiasing framework to tackle social biases, aiming to reduce the objectionable dependence between LLMs’ decisions and the social information in the input.
+Our framework introduces a novel perspective to identify how social information can affect an LLM’s decision through different causal pathways.
+Leveraging these causal insights, we outline principled prompting strategies that regulate these pathways through selection mechanisms.
+This framework not only unifies existing prompting-based debiasing techniques, but also opens up new directions for reducing bias by encouraging the model to prioritize fact-based reasoning over reliance on biased social cues.
+We validate our framework through extensive experiments on real-world datasets across multiple domains, demonstrating its effectiveness in debiasing LLM decisions, even with only black-box access to the model.
+Prompting Fairness: Integrating Causality to Debias Large Language Models
+11footnotetext: Equal contribution. Correspondence to: Jingling Li <jinglingli1024@gmail.com> and Zeyu Tang <zeyutang@cmu.edu>. The work was done while JL and YL were working at ByteDance Research.
 1 Introduction
-Large Language Models (LLMs), which are trained on massive text corpora, have been found to exhibit concerning levels of social biases (Sheng et al., 2019; Gonen & Goldberg, 2019; Schick et al., 2021; Bender et al., 2021; Dodge et al., 2021).
+Large language models (LLMs) trained on massive text corpora have been found to exhibit concerning levels of social biases (Sheng et al., 2019; Gonen & Goldberg, 2019; Schick et al., 2021; Bender et al., 2021; Dodge et al., 2021).
 The unchecked biases can potentially perpetuate and amplify societal inequities, leading to unfair or even unethical outcomes.
@@
 Many debiasing approaches have been proposed to tackle this issue, for instance, direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023).
-For various reasons such as security and business interests, many of the most capable LLMs are closed-sourced, for instance, GPT-4 (Achiam et al., 2023), Gemini (Anil et al., 2023), Claude 2 (Anthropic, 2023), where the general public do not have access to models’ internal structures or parameters.
-Thus, prompting-based techniques largely become the only viable option to mitigate bias when using closed-sourced LLMs.
-Figure 1: Different reasoning to answer a coreference question.
-A biased answer may be due to the use of a gender shortcut, while a bias-free answer is made by considering proper world knowledge given the circumstances.
-[^2] The use of demographic information does not necessarily indicate the reasoning is biased: sometimes certain demographic information (e.g., gender) should be considered for situations such as making medical decisions.
-In this work, we focus on prompting techniques to steer LLMs towards unbiased responses.
-To study this, we notice that obtaining unbiased responses essentially boils down to the process of selecting proper pieces from the model’s internal representations and knowledge to make the decision.
-Let us consider a simple task of resolving the coreference of a gender pronoun in a given sentence.
-A model may output a biased answer by using a gender shortcut potentially learned from the imbalanced representations in the training data.
-For example, in Figure 2(a), the model may associate the pronoun “he” with “physician” rather than “secretary” due to social biases reflected in the training data – the gender distributions are disproportional for different occupations.
-We denote such reasoning process as biased reasoning, where the model selects internal representations in an improper way as this specific demographic information (gender) is unrelated to solving the coreference task.
-We provide more detailed discussions on selection mechanisms in Section 3.
-Figure 2: Different prompting-based methods guided by one or both intuitions: discouraging ([-]) biased reasoning, and encouraging ([+]) bias-free reasoning.
-Most of the existing prompting-based debiasing methods focus on discouraging biased reasoning, for instance, using explicit prompts to avoid using biased associations or prohibiting the utilization of gender information when making the decision.
-While these approaches help to a certain extent, another overlooked prompting-based strategy is to encourage bias-free reasoning by only selecting proper pieces related to the nature of the question.
-For example, Figure 2(b) demonstrates one possible way of conducting bias-free reasoning by comparing the likelihood of two situations occurring in real life and drawing on the more plausible situation to infer the coreference resolution.
-We reveal the essential role of selection mechanisms in the interplay between LLM’s internal reasoning process and different strategies for designing the external prompts.
-We first construct a detailed causal model on the data generation process of the training corpus to identify ways biases may be smuggled in the pretraining phase.
-Then we construct a causal model of the LLM’s potential reasoning process and connect the two causal models by analyzing how different input prompts could modulate the LLM’s output through selection mechanisms.
-Building upon the above causal understandings, we introduce a causality-guided framework to debias LLMs, underpinned by the two intuitions: (1) discouraging biased reasoning  and (2) encouraging bias-free reasoning.
-Current prompting-based debiasing methods can also be viewed as instantiating one or both of the intuitions, as illustrated in Figure 2.
-Using the causality-guided debiasing framework, we propose prompt-design strategies and further conduct a variety of empirical studies to analyze their effectiveness.
-We find that prompts employing both intuitions significantly outperform existing prompting-based debiasing methods.
-The strong empirical results clearly demonstrate the effectiveness of the causality-guided debiasing framework, even when we only have black-box accesses.
+For various reasons such as security and business interests, many of the most capable LLMs are closed-source models, for instance, GPT-4 (OpenAI, 2023), Gemini (Anil et al., 2023), Claude (Anthropic, 2024), where the general public do not have access to models’ internal structures or parameters.
+Therefore, prompting-based techniques largely become the only viable option to mitigate bias when using closed-source LLMs.
+In this work, we focus on prompting techniques that mitigate bias in LLMs’ decisions.
+We observe that unbiased decision-making in an LLM essentially hinges on the selection and utilization of appropriate dependence patterns within its internal representations or knowledge.
+Let us consider a simple coreference resolution task on gender pronouns in a given sentence.
+A model may output a biased answer via a gender shortcut potentially learned from the unbalanced training data.
+For example, in Figure 1(a), the model may associate the pronoun “she” with “teacher” rather than “driver” due to social biases reflected in the training data–the gender distributions are disproportional for different occupations.
+We denote such reasoning process as biased reasoning, where the model selects and utilizes inappropriate dependence patterns involving specific social information (gender in this example) that should actually be irrelevant to solving the coreference task.
+(a) Different reasoning to a coreference question
+(b) A systematic way to promote fact-based reasoning
+Figure 1: Panel (a): A biased answer may be due to the use of a gender shortcut, while a fact-based answer is made by considering proper world knowledge given the circumstances.
+Panel (b): we describe how to systematically generate prompts that encourage fact-based reasoning.
+We would like to note that using social category information does not necessarily indicate that the reasoning is biased: sometimes, certain social category information should be considered, e.g., gender in medical treatments.
+We call such dependence neutral dependence, while in this work, we focus on objectionable/problematic ones (Tang et al., 2024).
+Most existing prompting-based debiasing methods focus on discouraging biased reasoning, for instance, using explicit instructions to avoid the usage of biased association when making the decision.
+While these approaches help to a certain extent, another overlooked debiasing strategy is to encourage fact-based reasoning by only selecting proper dependence patterns related to the nature of the question.
+For example, in Figure 1(a), we ask the model to first compare the likelihood of two situations occurring in real life and draw on the more likely situation to infer the coreference resolution.
+In Figure 1(b), we also provide a systematic approach to encourage fact-based reasoning.
+We formalize the above intuitions into a causality-guided debiasing framework, offering a novel perspective on how social information influences an LLM’s decision-making through different causal pathways.
+Within this proposed framework, we demonstrate how selection mechanisms can introduce bias into the training data while also help mitigate bias in LLMs’ decisions.
+Specifically, we present principled prompting-based strategies that regulate different causal pathways via employing selection mechanisms, and these strategies can be applied to (1) discourage biased reasoning, and (2) encourage fact-based reasoning.
+Current prompting-based debiasing methods can be viewed as instantiating one or both of these goals.
+We conduct extensive empirical studies on real-world datasets across multiple domains, and our proposed strategies significantly outperform existing prompting-based debiasing methods.
+The strong empirical results clearly demonstrate the effectiveness of our causality-guided debiasing framework, even when we only have black-box accesses.
 Our contributions can be summarized as follows:
-  • We construct detailed causal modelings for both the underlying data-generating process of the training corpus and the LLM reasoning process.
-We note that selection mechanisms play an essential role in identifying how the LLM’s output could be modulated by different prompts.
-  • We formulate a causality-guided debiasing framework.
-Through causal understandings of data-generating processes, we identify conditions that strategies for designing prompts should satisfy to debias in a principled and effective way.
-The framework also unifies existing prompting-based debias methods.
-  • Using these strategies, we show strong empirical results on debiasing various social biases across different LLMs, demonstrating the clear benefit of our framework.
+  • We develop a causality-guided debiasing framework that provides a novel perspective on how social information can influence LLMs’ decisions through different causal pathways.
+  • Using the proposed framework, we present principled prompting-based debiasing strategies that utilize selection mechanisms to regulate the influence of bias.
+These strategies aim to discourage biased reasoning and encourage fact-based reasoning, offering a unified view of existing prompting-based methods and addressing the gap in promoting fact-based reasoning.
+  • We conduct extensive empirical studies and demonstrate the effectiveness of our framework in debiasing LLMs’ decisions across various social categories, highlighting both its theoretical foundations and practical significance in debiasing LLMs’ decisions.
 2 Preliminaries
-In this section, we provide a brief introduction to causal modeling and causal reasoning (Section 2.1).
-We also provide a motivating example to illustrate how the selection mechanism can reshape dependence patterns within the different data (Section 2.2).
-[^3] We discuss related works in detail in Appendix A.
+In this section, we introduce our problem setting, defining what we mean by an unbiased decision for large language models (Section 2.1).
+Then we provide a brief introduction to causal modeling and causal reasoning (Section 2.2).
+We also provide a motivating example to illustrate how the selection mechanism can reshape dependence patterns among different variables (Section 2.3).
+2.1 Problem Settings
+[…]
 2.2 A Brief Introduction to Causality
-For two random variables $X$ and $Y$, $X$ is a cause of $Y$ if there is a change in the distribution of $Y$ when we apply an intervention on $X$ while holding all other variables fixed (Spirtes et al., 1993; Pearl, 2009).
+For two random variables $X$ and $Y$, $X$ is a direct cause of $Y$ if there is a change in the distribution of $Y$ when we apply an intervention on $X$ while holding all other variables fixed (Spirtes et al., 1993; Pearl, 2009).
 We can represent causal relations among variables with a directed acyclic graph (DAG), where nodes represent variables, and edges represent direct causal relations between variables.
 We denote the direct causal relation between the ordered pair $(X,Y)$ by a directed edge $X\rightarrow Y$.
-Local causal modules in a DAG, which characterize the causal relations between the corresponding variable and its direct causes, do not interfere with each other because of causal modularity.
-This property is also known as the exogeneity (Engle et al., 1983), or the independence of causal mechanism (Peters et al., 2017), resulting directly from the causal Markov condition for the DAG (Spirtes et al., 1993; Pearl, 2009).
-[^4] There are additional classes of graphs considered in causality literature, for example, directed cyclic graphs (DCGs) (Spirtes, 1995), ancestral graphs (Richardson & Spirtes, 2002), and so on.
-In this paper, we consider causal processes that can be modeled by a DAG.
-Other graph classes are beyond the scope of our work.
 In the context of language processing, the definition of a variable representing text or tokens is relatively abstract compared to the statistical notion of a random variable in tabular data.
 Within the scope of this work, we use the terms “variable” and “node” interchangeably when the context permits clear understandings.
-2.2 A Motivating Example of Selection Mechanisms
+2.3 Remarks on Selection Mechanisms: An Illustrative Example
 Ideally, we would like samples to be drawn uniformly from the underlying population of interest.
-However, in practice, it is very common that the probability of including certain data points in the corpus depends on the characteristics of the data points themselves.
-Previous literature has investigated selection mechanisms from different perspectives, for instance, the influence of selection bias on statistical inference in economic and sociological studies (Heckman, 1979, 1990; Winship & Mare, 1992), causal discovery when there are selection variables and latent common causes (Spirtes et al., 1995; Zhang, 2008), the identification and estimation of functional causal models when selection exists (Zhang et al., 2016), the identifiability of causal effect in the presence of selection bias from the graphical condition perspective (Bareinboim & Pearl, 2012; Bareinboim & Tian, 2015; Correa et al., 2019) and from the potential outcome perspective (Hernán & Robins, 2020), and the identification of the existence of selection bias from observational data under certain functional assumptions (Kaltenpoth & Vreeken, 2023).
+However, in practice, the probability of including certain data points in the training corpus often depends on their characteristics where selection mechanisms are commonly involved.
+[^1] A thorough literature review on selection mechanisms can be found in Appendix A.1.
 Figure 2: An illustrative example of selection mechanisms.
-Let us consider an example in the context of medical care.
-As illustrated in Figure 3, observed variables $(X_{1},X_{2})$ denote diseases, and $(Y_{1},Y_{2})$ denote corresponding symptoms.
-Apart from them, there exist potential binary selection variables $S_{i}$’s ($i\in\{1,2,\ldots,5\}$), where $S_{i}=1$ denotes being selected.
-$X_{1}$ ($X_{2}$) is the direct and only cause of $Y_{1}$ ($Y_{2}$), and the two diseases are unrelated in the general population, i.e., when none of the $S_{i}$’s exists.
-We use solid edges to represent causal relations among observed variables and dashed edges for those pertaining to selection mechanisms.
+Consider an example in the context of medical care: as illustrated in Figure 2, the observed variables $(X_{1},X_{2})$ denote diseases and $(Y_{1},Y_{2})$ represent corresponding symptoms.
+There can exist potential binary selection variables $S_{i}$’s ($i\in\{1,2,\ldots,5\}$), where $S_{i}=1$ indicates the occurrence of that selection.
+$X_{1}$ ($X_{2}$) is the direct and only cause of $Y_{1}$ ($Y_{2}$), and the two diseases are uncorrelated in the general population, i.e., when none of the $S_{i}$’s exists.
+Solid edges in the figure denote causal relations among observed variables, while dashed edges represent the ones pertaining to selection mechanisms.
 For instance, $S_{2}$ is an example of outcome-dependent selection (Zhang et al., 2016), which can be selecting individuals with symptom $Y_{1}$ from the general population.
-The conditional probability $P(Y_{1}\mid X_{1},S_{2}=1)$ in the selected data typically differs from its counterpart $P(Y_{1}\mid X_{1})$ in general population (Zhang et al., 2016).
-As another example, $S_{4}$ can denote the selection mechanism of only considering the hospital in-patient data.
-This signifies the setting of the well-known Berkson’s Paradox (Berkson, 1946), where two unrelated diseases $(X_{1},X_{2})$ appear to be correlated in the hospital data, simply because the data only contains selected patients who have at least one symptoms in $(Y_{1},Y_{2})$.
-Figure 3 shows that the selection can be based solely on the cause (e.g., $S_{1}$), solely on the effect (e.g., $S_{2}$), or on both (e.g., $S_{3}$).
-Meanwhile, the selection can be based on variables that are not causally related in the general population (e.g., $S_{4}$ and $S_{5}$).
-Selection mechanisms can reshape dependence patterns among involved variables, and as a consequence, potentially change downstream outputs as well.
-We will see in Section 3 that such property of selection also applies to natural language processing (NLP) contexts.
+The conditional probability $P(Y_{1}\mid X_{1},S_{2}=1)$ in the selected data typically differs from its counterpart $P(Y_{1}\mid X_{1})$ in the general population (Zhang et al., 2016).
+As another example, $S_{4}$ demonstrates the selection solely based on the hospital in-patient data, a setting of the well-known Berkson’s Paradox (Berkson, 1946), where two unrelated diseases $(X_{1},X_{2})$ appear to be correlated in the hospital data, simply because the sample includes only admitted patients who have at least one symptoms in $(Y_{1},Y_{2})$.
+[…]
+3 Causality-Guided Debiasing Framework
+This section presents our causality-guided debiasing framework.
+Specifically, we provide a causal modeling on the data generating process of LLMs’ training corpora in Section 3.1, highlighting how bias can be introduced into the training data by selection mechanisms.
+In Section 3.2, we demonstrate how social category information can impact LLMs’ decisions via different causal pathways.
+Section 3.3 further reveals how selection mechanisms can regulate the information flow over these pathways by reshaping dependence patterns.
+The regulations over different causal pathways leads to prompting strategies with distinct objectives aimed at effectively mitigating biases in LLMs.
+Moreover, we prove that when these objectives are satisfied, we can completely remove the influence of social category information from the LLMs’ decisions, which allows the LLMs’ decision to be conditionally independent of the social information, given the presence of appropriate selection mechanisms.
 (a) Training data corpus
-(b) LLM reasoning process given prompt
-(c) LLM reasoning process (annotated copy)
-Figure 4: Causal graphs with respect to different data-generating processes.
-Panel (a) presents the underlying data-generating process of the training data corpus.
-Panels (b) – (c) present the reasoning process of the LLM.
-Since the conditions specified by the prompt are always conditioned upon when LLM generates the output, the internal reasoning is modulated by the external prompt.
-In panel (c) we include annotations to highlight causal pathways along which the information flow from demographic representation to LLM potential output is not regulated.
-3 Causality-Guided Debiasing Framework
-We propose our debiasing framework guided by causal understandings of involved data-generating processes.
-In Section 3.1, we present detailed causal models of both the underlying data-generating process for the training corpus and the LLMs’ reasoning process.
-We reveal that the LLM’s reasoning process is essentially an interplay between its internal representations and conditions specified by the external prompt designs, in which selection mechanisms play a key role.
-In Section 3.2, we identify different sets of conditions for prompt designs, which lead to prompting strategies that are not only intuitive but also theoretically grounded.
-3.1 Detailed Causal Modeling of Processes
-We present the causal modeling of the underlying data-generating process of the training data corpus (Section 3.1.1) and the reasoning process of LLMs (Section 3.1.2).
-3.1.1 Underlying Data Generating Process of Text in Training Data Corpus
-The training data corpus often reflects historical discrimination, and selection mechanisms are commonly involved.
-For instance, gender stereotype in occupations arises not due to the existence of a direct causal relation or a common cause between gender and occupation, but due to an underlying selection mechanism.
-Specifically, in the training data, among all possible combinations between gender (e.g., male and female) and occupation (e.g., CEO and secretary), there is a tendency to associate CEO more often with male, and secretary with female.
-This phenomenon occurs because the training data is a subset selected from an imaginary data corpus that is ideally diverse and comprehensive.
-More generally, in Figure 4(a) we model causal relations in the underlying data-generating process of training data corpus (e.g., text scraped from the internet).
-Other than demographic information (e.g., race, gender, age), the causal graph contains additional variables of interest.
-We use “scenario” to represent the practical situation or context that sets the background of the text (e.g., medical care, hiring).
-We use “entity” to denote participants or stakeholders involved in the scenario, for instance, a patient may be an entity in the medical care scenario and a secretary may be an entity in the hiring scenario.
-The selection variable $S$ explicitly models the association between demographic information and entity, which is recognized as a major type of stereotype in NLP (Sweeney, 2013; Bolukbasi et al., 2016; Zhao et al., 2018; Tamkin et al., 2023).
-There are different types of texts in Figure 4(a).
-“Demographic-agnostic fact” denotes the text that does not explicitly contain demographic information.
-“Derivational text based on fact” denotes the text derived from fact, e.g., inference according to definition, fact-check Q&A, and restatement without altering factual contents.
-“Demographic-aware text” denotes the text where demographic information appears explicitly.
-3.1.2 Reasoning Process of LLMs
-Section 3.1.1 introduces causal modeling of the underlying data-generating process to characterize how discrimination is instantiated in LLMs’ training data corpus.
-The goal and expectation of the training is for LLM models to capture dependence patterns in the data corpus, and to extract and internalize the related knowledge.
-Therefore, we assume that the internal reasoning process of LLMs shares similarities with the underlying data-generating process of data corpus.
-Under this mild assumption, we formulate the reasoning process of how LLMs generate outputs based on input prompts in Figure 4(b).
-In particular, we reveal the interplay between internal representations and conditions specified by external inputs, and more importantly, how LLM outputs are shaped and modulated by different prompt designs through selection mechanisms.
-In Figure 4(b), we use dotted contours to distinguish LLMs’ internal representations from the actual external information in the data corpus, e.g., the contrast between “demographic representation” in Figure 4(b) and “demographic information” in Figure 4(a).
+(b) How input prompts modulate LLM decisions
+Figure 3: Causal graphs for different data generating processes.
+We use double-stroke contours to indicate selection variables, solid edges to represent causal relations among observed variables or internal representations, and dashed edges for those pertaining to selection mechanisms.
+Panel (a) presents the underlying data generating process of the training data corpus.
+Panel (b) presents a causal perspective on how LLM’s decisions are related to internal representations and modulated by external prompts.
+We highlight in light coral to denote pathways along which social category information can influence the LLM’s decision.
+3.1 Underlying Data Generating Process of LLMs’ Training Corpora
+The training data of LLMs often reflects historical discrimination or stereotypes, and selection mechanisms are commonly involved.
+For instance, gender stereotype in occupations arises not due to the existence of a direct causal relation or a common cause between gender and occupation, but rather from an underlying selection mechanism: in the training data, among all possible combinations between gender (e.g., male and female) and occupation (e.g., CEO and secretary), there is a tendency to associate CEO more often with male, and secretary with female due to historical or societal biases.
+This phenomenon may occur when the training data is a subset selected from an imaginary data corpora that is ideally diverse and comprehensive, where there is no historical discrimination or stereotype when people generate text content.
+Therefore, the selection mechanism between gender and occupations introduces bias into the training data, which was fed to LLMs during the pretraining stage.
+More generally, in Figure 3(a), we provide a causal graph to denote the underlying data generating process of training corpora (e.g., text scraped from the internet).
+The causal graph contains additional variables of interest besides information related to a social category (e.g., age.
+gender, and race): “scenario” represents the practical situation or context where a decision is needed (e.g., medical care, hiring); “entity” denotes the participants or stakeholders involved in the scenario (e.g., a patient in the medical care scenario, or a secretary in the hiring scenario); and the selection variable $S$ explicitly models the associations of social categories with particular entities and scenarios, which are recognized as a major type of stereotypes in NLP (Sweeney, 2013; Bolukbasi et al., 2016; Zhao et al., 2018; Tamkin et al., 2023).
+There are also different types of texts in Figure 3(a): “social-agnostic fact” denotes the text that does not explicitly contain information related to a social category of interest (e.g., base question in Figure 1(b)); “social-salient text” denotes the text where information related to the social category appears explicitly (e.g., original question in Figure 1(b)); and “derivational text based on fact” denotes the text or decision derived from factual evidence (e.g., inference according to fact-based reasoning).
+3.2 How Social Category Information Can Impact LLM’s Decisions
+The causal modeling of the underlying data generating process in Section 3.1 characterizes how bias could be introduced into LLMs’ training data.
+In this subsection, we explore how LLM outputs can be potentially shaped and modulated by different prompt designs through selection mechanisms, as presented in Figure 3(b).
+We highlight certain edges in light coral (accompanied by annotations marked with circled red numerals) to denote unregulated information flow from social category representations to LLM’s output, which potentially leads to biases in the output decisions.
+We use dotted contours to distinguish LLMs’ internal representations from the actual external information in the data corpora, e.g., the contrast between “social category” in Figure 3(a) and “social category representation” in Figure 3(b).
 Notice that the internal nodes are not directly observable or accessible.
-We use double-stroke contours to indicate selection mechanisms, directed edges to represent direct causal relations, and dashed edges with hollow arrowheads to denote selection mechanisms.
-A “prompt” serves as an external input to LLMs but not as a direct cause of internal representations.
-This is because the internal knowledge and representations exist beforehand, making them irrelevant to whether a specific prompt is provided.
-The prompt also does not act as an indicator for causal interventions.
-Because internal nodes are not directly observable or accessible, one cannot set them to certain values via hard interventions (Spirtes et al., 1993; Pearl, 2009; Peters et al., 2017; Hernán & Robins, 2020), or change the functional behavior of causal modules via soft interventions (Eberhardt & Scheines, 2007; Huang et al., 2020; Correa & Bareinboim, 2020).
-However, the prompt directly changes the selection variable “prompt properly considered” (PPC) through its designs.
-If the LLM model is well-trained and well-aligned, we can expect that PPC is always conditioned upon when the model produces the output, and the “LLM potential output” will become the actual output we obtain from the model.
-Just like the motivating example where selection can reshape dependence patterns among involved variables (Section 2.2), the prompt can influence the LLM reasoning process in significant ways by specifying conditions in selection mechanisms.
-3.1.3 A Remark on the Modeled Processes
-Figure 4(a) and Figure 4(b) each have their generating process of interest and distinct emphases.
-Although not pertaining to LLM reasoning itself, the underlying generating process modeled in Figure 4(a) provides hints on local causal modules of interest that prompt designs can specifically attend to for debiasing purposes.
-When we consider these hints with reference to the LLM reasoning process modeled in Figure 4(b), we can identify debiasing strategies that are both intuitive and theoretically grounded, as we will see in more detail in Section 3.2.
-The two detailed causal models complement each other, both of which are essential for understanding the source of bias, and furthermore, effectively debiasing LLMs.
-3.2 Debiasing Guided by Causal Understandings
-We present our prompt-based LLM debiasing framework guided by causal understandings of the related data-generating processes.
-Our core idea is to formulate conditions that should be specified in the prompt design, such that through the influence of selection mechanisms in the LLM reasoning process, one can effectively debias LLM outputs.
-Figure 4(c) presents an annotated version of the LLM reasoning process (Section 3.1.2).
-Based on the understanding of the underlying generating process of training data corpus, and the mild assumption that the trained LLM captures the dependence patterns in training data, we highlight certain edges in light coral (accompanied by annotations marked with circled red numerals) in Figure 4(c) to denote unregulated information flow from demographic representations to LLM outputs in the internal reasoning process.
-We use circled blue Roman numerals to denote selection mechanisms that can be specified by the external input prompt (Section 3.1.2), and circled red Roman numerals to represent those corresponding to historical discrimination (Section 3.1.1).
-According to the LLM reasoning process presented in Figure 4(c), we consider the information flows from the demographic representation (upstream node) to the LLM potential output (downstream node).
-We present additional conditions and constraints the prompt designs should specify for debiasing purposes.
-To clearly present the intuitions and theoretical groundings of our causality-guided framework, we identify three prompting strategies for debiasing LLMs, each of which serves as a solid starting point.
-Strategy I (Nudge Towards Demographic-Agnostic Fact).
-The intuition behind is to nudge LLMs towards utilizing demographic-agnostic fact when generating the output:
-  1. Condition I
-    The internal representations for demographic-agnostic fact and demographic information should be conditionally independent in the presence of PPC and existing selection $S$, i.e.,
+For the selection mechanisms in Figure 3(b), we use circled red Roman numerals to represent ones corresponding to historical discrimination (Section 3.1), and we use circled blue Roman numerals to denote the ones that can be specified by the external input prompt.
+In Figure 3(b), a “prompt” serves as an external input to LLM rather than a direct cause of changes in internal representations, because the internal knowledge and representations exist a priori, and whether a specific prompt is provided does not alter their existence.
+The prompt also does not act as an indicator for causal interventions, since internal nodes are not directly observable or accessible, and one cannot set them to certain values via hard interventions (Spirtes et al., 1993; Pearl, 2009; Peters et al., 2017; Hernán & Robins, 2020), or change their functional behavior via soft interventions (Eberhardt & Scheines, 2007; Huang et al., 2020; Correa & Bareinboim, 2020).
+However, the input prompt directly changes the selection variable “prompt properly considered” (PPC), a concept which we further examine in Section 3.3.
+If the LLM model is well-trained and well-aligned, we can expect that the model will always condition on PPC when producing a “LLM’s potential decision.”
+3.3 Causality-Guided Debiasing Strategies
+As illustrated in Figure 3(b), social category information can influence the LLM’s potential decisions via two important nodes: the “social-agnostic fact representation” node and the “social-salient text representation” node.
+Our core idea is to formulate objectives that regulate the information flow through these two nodes.
+By designing prompts that satisfy these objectives, we can effectively debias LLM decisions through selection mechanisms.
+Regulating the information flow through the “social-agnostic fact representation” node leads to prompt-design strategies aimed at encouraging fact-based reasoning (Strategy I and Strategy II), and regulating the flow through the “social-salient text representation” node results in strategies that discourage biased reasoning (Strategy III).
+Strategy I (Nudge Towards Social-Agnostic Fact).
+Strategy I steers LLMs towards relying more on the social-agnostic fact when making decisions.
+The objective is to utilize the selection mechanism over internal representations of social category (connected via edge i) and social-agnostic fact (connected via edge v), aiming to make them conditionally independent in the presence of PPC and the existing selection $S$:
 […]
-Strategy I introduces Condition I to specify the selection mechanism over internal representations of demographic information and demographic-agnostic fact, denoted by edges i and v in Figure 4(c).
-An example prompt employing Strategy I can be: “Considering the fact that the sentence ‘The physician hired the secretary because the secretary is highly recommended’ is practically more viable than the sentence ‘The physician hired the secretary because the physician is highly recommended’, who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
-Strategy II (Counteract Existing Selection Bias).
-The intuition behind this strategy is to directly counteract the effect of existing historical discriminations:
-  1. Condition II.1
-    The internal representations for demographic information and entity should be conditionally independent in the presence of PPC and existing selection $S$, i.e.,
+Strategy I focuses on social-agnostic fact, which is a direct parent of LLM potential decision.
+The objective is to encourage LLMs to utilize social-agnostic fact, while regulating its dependence on social category through the selection mechanism involving these two (denoted by edges i and v).
+The ideal outcome is that the information flow through social-agnostic fact to downstream nodes (along edges 4 and 5) is regulated (see Figure 3(d)).
+An example prompt employing Strategy I can be: “Considering the fact that the sentence ‘the driver offered a ride to the teacher but the driver had an accident on the way’ is more likely to happen in real life than the other sentence ‘the driver offered a ride to the teacher but the teacher had an accident on the way,’ who does ‘she’ refer to in the sentence ‘the driver offered a ride to the teacher but she had an accident on the way’?”
+Strategy II (Alternative Fact Counteracts Existing Selection Bias).
+Strategy II directly counteracts existing bias instantiated by the existing selection $S$ (through edges { vi, vii, viii}).
+The objective is to use selection mechanisms over internal representations of social categories, entities, scenarios (through edges { i, ii, iii}) to enforce conditional independence between entity and social category, and between scenario and social category, in the presence of PPC and the selection $S$:
 […]
-  2. Condition II.2
-    No new association between internal representations of demographic information and demographic-agnostic fact is introduced,
+Strategy II aims to counteract the existing selection bias from $S$ by introducing corresponding selection mechanisms that simultaneously operate over internal representations for social category (by PPC via edge i and by $S$ via edge vi), entity (by PPC via edge ii and by $S$ via edge vii), and scenario (by PPC via edge iii and by $S$ via edge viii).
+The ideal outcome is that the downstream node of entity and scenario representations, i.e., social-agnostic fact, does not inherit dependence on the social category from its direct parents (along edges 2 and 3; see Figure 3(e)).
+An example prompt employing Strategy II can be: “ Assume male and female are equally represented in drivers and in teachers.
+Who does ‘she’ refer to in the sentence ‘the driver offered a ride to the teacher but she had an accident on the way’?” despite that the actual representation may be highly uneven in the historical data used in pretraining.
+Strategy III (Nudge Away from Social-Salient Text).
+Strategy III steers LLMs away from utilizing social-salient information when making decisions.
+The objective is to utilize the selection mechanism over internal representations of social categories (via edge i) and social-salient text (via edge iv) to make them conditionally independent in the presence of PPC and the existing selection $S$:
 […]
-In Strategy II, Condition II.1 and Condition II.2 serve different purposes.
-Condition II.1 aims to counteract existing bias instantiated by selection $S$ (edges vi and vii) by constraining marginal dependence between representations for demographic information and entity (edges i and $i$).
-Condition II.2 acts as a safeguard, making sure no new bias is introduced in the causal downstream of the entity so that the above counteraction effectively proceeds to the final output.
-An example prompt employing Strategy II can be: “The physician can be either male or female, and the secretary can also be either male or female.
-Who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
-Strategy III (Nudge Away from Demographic-Aware Text).
-The intuition behind this strategy is to nudge LLMs away from utilizing demographic-aware text to generate outputs:
-  1. Condition III
-    The internal representations for demographic-aware text and demographic information should be conditionally independent in the presence of PPC and existing selection $S$,
+Strategy III focuses on social-salient text, which is a direct parent of LLM potential decision.
+Since social-salient text inherently contains information about social category, Strategy III essentially discourage the use of social-salient text when forming decisions.
+This is achieved by regulating the information flow between social category and LLM potential decision through edges 1 and 6 (see Figure 3(f)).
+An example prompt employing Strategy III can be: “Do not answer the question using gender information.
+Who does ‘she’ refer to in the sentence ‘the driver offered a ride to the teacher but she had an accident on the way’?”
+Remarks on the Three Strategies.
+While each of the three strategies is individually effective to some extent, none is perfect in isolation.333 We provide additional illustrations in Appendix B.
+For instance, although Strategy I steers LLMs towards using social-agnostic facts, it does not explicitly prevent LLMs from using social-salient representations in making decisions.
+Consequently, the social category can still influence the output decision through an unregulated path involving edges 1 and 6.
+Similarly, while Strategy II aims to regulate information related to social category representation that flows from representations of entity and scenario to downstream along edges { 2, 3}, there is no explicit constraint involving the information flow along edges 1 and 6, or along edges 4 and 5, which leaves space for bias to sneak in.
+In contrast to Strategy I, which encourages LLMs to focus on social-agnostic facts, Strategy III prevents LLMs from relying on social-salient text during reasoning.
+Although edge 1 is explicitly regulated by the selection mechanism denoted by edges i and iv, the information flow along edge 5 can still result in the association between social category and the output decision.
+This is because Strategy III does not involve objectives to change the dependence patterns between social category and social-agnostic fact.
+As seen from the objective specified in Strategy II, the absence of explicit social category information in a social-agnostic fact does not guarantee the absence of dependence between the two, since direct parents (entity and scenario representations) of social-agnostic fact may still be dependent with the social category if without explicit constraints.
+Debiasing is better realized when these strategies are combined to address social bias in LLMs more comprehensively:
+Theorem 3.1 (Comprehensive Debiasing When Combining All Three Strategies).
+If conditions and constraints specified in Equations (1), (2), and (3) are simultaneously satisfied, the LLM’s decision $Y$ is independent from the social category $A$ in the presence of PPC and existing selection $S$:
 […]
-Strategy III utilizes Condition III to specify the selection mechanism over internal representations of demographic information and demographic-aware text (edges i and iv) to regulate the information flow along the edge 1.
-An example prompt employing Strategy III can be: “Do not use gender information to answer the question. Who does ‘he’ refer to in ‘The physician hired the secretary because he is highly recommended’?”
-Remarks on Three Strategies
-The three strategies offer certain effectiveness individually but are not perfect on their own.
-[^5] We provide additional illustrations in Appendix B.
-While Strategy I nudges LLMs towards utilizing demographic-agnostic facts, it does not explicitly prevent LLMs from using demographic-aware text representations to generate the output, and the demographic information can potentially be associated to the output through an unregulated path containing edges 1 and 5.
-Similarly, while Strategy II aims to regulate information flows along edges { 2, 3, 4, vi, vii}, there is no explicit constraint involving edges 1 and 5, which leaves space for bias to sneak in during the reasoning process.
-Compared to Strategy I that pushes LLMs to focus on demographic-agnostic facts, Strategy III prevents LLMs from referring to demographic-aware text during reasoning.
-While edge 1 is explicitly regulated by the selection mechanism specified by Condition III, the information flow along edge 5 can still result in the association between demographic information and the output.
-This is because Strategy III does not involve conditions or constraints over the dependence patterns among variables including demographic information and demographic-agnostic fact.
-As we can see from the conditions specified in Strategy II, no explicit appearance of demographic information in a demographic-agnostic fact does not guarantee the non-existence of dependence between the two.
-We would like to note that debiasing is better realized when the strategies are combined as they can address social bias in LLMs more comprehensively.
+Remark 3.2.
+The PPC here is the selection node representing multiple selection mechanisms simultaneously, with the objectives in all three prompting strategies fully achieved.444 We implicitly utilize the mild assumption that a well-trained and well-aligned LLM captures the dependence pattern in the training data and that such a pattern is internalized and utilized during reasoning.
+Though it may be hard to realize full achievements in practice, we show in our experiments that even if we cannot fully satisfy any of the three objectives, we can still design prompts following the proposed strategies to effectively debias LLMs’ decisions.
+Proof.
+There are two direct parents of LLM’s decision $Y$.
+To enforce the independence between $Y$ and the social category $A$, a sufficient condition is that both direct parents of $Y$ are conditionally independent of $A$, in the presence of proper PPC and the existing selection $S$.
+For the social-agnostic fact, considering its direct parents–representations for entity and scenario–the conditional independence can be indirectly enforced through Equation (2), and further directly guaranteed by Equation (1).
+For social-salient text, the conditional independence can be enforced by Equation (3).
+Therefore, Equations (1), (2), and (3) altogether removes the (objectionable) dependence between social category information and LLM decisions, as described in Equation (4).
+∎
+Theorem 3.1 indicates that when jointly employing Strategies I, II, and III, we can successfully achieve the debiasing of LLM decisions.
+From a purely technical standpoint, Equation (1) and Equation (2) might seem redundant, as both of them aim to enforce the conditional independence between social category and social-agnostic fact given PPC and $S$.
+However, because of the definition of a variable in the language context is relatively abstract, it is favorable to have empirical strategies that can debias LLM decisions from different perspectives.
+In Section 4, we empirically demonstrate the advantage of our strategies, when utilized individually and in combination.
+4 Results and Analyses
+In this section, we measure how (1) encouraging fact-based reasoning (Strategies I and II), and (2) discouraging biased reasoning (Strategy III), effectively steer LLMs towards making unbiased decision across various social dimensions.
+We conduct extensive experiments on three widely utilized benchmarks that evaluate language models’ decision bias: WinoBias by Zhao et al.
+(2018) (Section 4.1), the Bias Benchmark for QA (BBQ) by Parrish et al.
+(2021) (Section 4.2), and Discrim-Eval by Tamkin et al.
+(2023) (Appendix D.2).
+On all three benchmarks, our framework leads to strategies effectively reduce LLMs’ decision biases across different social categories.
+4.1 Gender Bias: WinoBias
+The WinoBias dataset (Zhao et al., 2018) evaluates how likely models assign stereotypical gender pronouns to occupations in coreference resolution tasks.
+The sentences in WinoBias are designed to be structurally parallel but differ in gender pronouns.
+It contains two sets of sentences: pro sentences with the pro-stereotypical gender pronouns (e.g., nurses as she, engineers as he), and anti sentences with anti-stereotypical gender pronouns (e.g., nurses as he, engineers as she).
+The dataset also has two types of tasks with different levels of difficulties: coreference decisions in Type I task are challenging and mandate world knowledge to reason about given circumstances, whereas Type II task can be resolved using only syntactic information.
+4.1.1 Experimental Settings (WinoBias)
+For each sentence in WinoBias, we define the original question to be “Who does {gender pronoun} refer to in the sentence ‘{original sentence}’?”, and we measure the performance of four large language models: GPT-3, GPT-3.5, GPT-4, and Claude 2 across the above two types of coreference tasks (Type I and Type II).
+Baseline Approaches
+There are a few existing works employing prompting techniques to mitigate the bias in LLMs. We consider these three baselines:
+  • Default: asks the original question directly without any additional prompts.
+  • ICL with contrastive examples: provides pairs of contrastive examples followed by the original question.
+The pair of questions are contrastive as the pronouns are different but the answer remains unchanged.
+In WinoBias, we use the same 16 ICL examples as in Si et al.
+(2022).
+Oba et al.
+(2023) also used similar counterfactual preambles to suppress bias in LLMs.
+  • Zero-shot COT (Kojima et al., 2022): asks the model to also “think step by step” after the original question.
+Shaikh et al.
+(2022); Ganguli et al.
+(2023) designed similar COT approaches to mitigate bias in LLMs.
+Evaluation Metrics
+We measure the accuracy of LLMs on pro and anti sentences, and the gap between the two indicates the level of gender bias exhibited by the models (smaller gaps are better).
+Dual Directional Prompting
+We propose Dual Directional Prompting (DDP) as one way of combining different prompting strategies.
+Following Figure 1, we construct the base question by first creating two gender-agnostic sentences via replacing the gender pronoun with the two occupations appeared in the original sentence, and then asking the model which sentence is more likely to happen in real life.
+This base question does not contain any gender-related information.
+We then prepend the model’s answer (to the base question) before asking the original question.
+This allows us to distill LLM’s non-gender-related world knowledge and nudge it to explicitly reason with this fact during coreference resolution (Strategy I).
+On top of Strategy I, we also add the prompt that both occupations are equally likely to be male or female to in order to counteract existing selection bias (Strategy II).
+Note that the generation of base questions can be done by regular expression or one additional LLM query.
+For efficiency, we can also query a smaller LLM to generate the base question given the original one and the related social category.
+One caveat of utilizing LLM’s world knowledge is that the performance of Dual Directional Prompting will increase as the capabilities of the LLM grow, since better world knowledge will further help LLM answer both factual and original questions.
 Table 1: Performance comparison of various debiasing methods on WinoBias.
-We show that combining encouraging bias-free reasoning and discouraging biased reasoning together (Reduce + Fact) significantly alleviates the gender bias for both coreference tasks: Type I coreference resolution requires world knowledge, and Type II coreference resolution can be resolved using only syntactic information.
-Pro in the table stands for coreference with pro-stereotypical pronouns, and Anti stands for coreference with anti-stereotypical pronouns.
+We show that combining encouraging fact-based reasoning and discouraging biased reasoning together (i.e., DDP) significantly alleviates the gender bias for both coreference tasks.
+“Pro” in the table stands for coreference with pro-stereotypical pronouns, and “Anti” stands for coreference with anti-stereotypical pronouns.
 A smaller gap indicates less bias.
-Accuracy (%) GPT 3 GPT 3.5 Claude 2 GPT 4
-Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$ Anti Pro $\text{Gap}_{\downarrow}$
+Accuracy (%) GPT-3 GPT-3.5 Claude 2 GPT-4
+[…]
 Type I
@@
 COT (zero shot) 41.79 75.85 34.06 62.14 90.64 28.49 70.56 91.59 21.03 84.40 95.66 11.26
-ICL 46.81 94.57 47.76 45.18 92.81 47.63 73.68 92.27 18.59 88.87 98.10 9.23
-Reduce + Fact 73.27 73.95 0.68 72.73 84.67 11.94 74.08 75.17 1.09 94.57 96.74
-Type II    Default 85.01 97.97 12.96 94.28 98.98 4.70 93.77 97.46 3.68 97.59 99.49 1.91 COT (zero shot) 69.12 86.79 17.66 93.90 98.35 4.45 95.30 99.62 4.32 97.97 99.62 1.65 ICL 94.41 99.62 5.21 95.43 98.98 3.56 94.03 97.84 3.81 98.35 99.87 1.52 Reduce + Fact 83.23 84.24 1.02 92.12 94.41 2.29 82.34 85.26 2.92 99.62 99.75 0.13
-4 Results and Analyses
-In this section, we demonstrate how the two types of approaches, i.e., (1) encouraging bias-free reasoning (Strategy I), and (2) discouraging biased reasoning (Strategy II and Strategy III), effectively steer LLMs towards unbiased response on various aspects of social bias (e.g., gender, race, and age).
-We present our experimental settings and results on the WinoBias data set (Zhao et al., 2018) in Section 4.1, and these on the Discrim-Eval data set (Tamkin et al., 2023) in Section 4.2.
-4.1 Gender Bias: WinoBias
-The WinoBias data set (Zhao et al., 2018) evaluates how likely models will assign stereotypical gender pronouns to occupations under coreference resolution tasks.
-The sentences in WinoBias are designed to be structurally parallel but differ in gender pronouns.
-It contains two sets of sentences: pro sentences with the pro-stereotypical gender pronouns (e.g., nurses as she, engineers as he), and anti sentences with anti-stereotypical gender pronouns (e.g., nurses as he, engineers as she).
-The data set also has two types of tasks with different levels of difficulties: coreference decisions in Type I task are challenging and must be made using world knowledge about given circumstances, whereas Type II task can be resolved using only syntactic information.
-4.1.1 Experimental Settings (WinoBias)
-For each sentence in WinoBias, we define the original question to be “Who does [gender pronoun] refer to in the sentence ‘[original sentence]’?”, and we measure the performance of four large language models: GPT-3, GPT-3.5, GPT-4, and Claude 2) across the above two types of coreference tasks (Type I and Type II).
-Baseline Approaches
-There are a few existing works employing prompting techniques to mitigate the bias in LLMs. We consider these three baselines:
-  • Default: asks the original question directly without any additional prompts.
-  • ICL with contrastive examples: provides contrastive examples followed by original question (Figure 2 middle).
-The examples are considered to be contrastive as the answer remains unchanged with different pronouns.
-In WinoBias, we use the same 16 ICL examples as in Si et al.
-(2022).
-  • Zero-shot COT (Kojima et al., 2022): asks the model to also “think step by step” after the original question.
-Shaikh et al.
-(2022); Ganguli et al.
-(2023) design similar COT approaches to mitigate bias in LLMs.
-Evaluation Metrics
-We measure the performances of LLMs on pro and anti sentences in terms of accuracy, and the gap between the two indicates the level of gender bias exhibited by the models (a smaller gap is better).
-Our Method: Reduce + Fact
-We propose Reduce + Fact as one way of encouraging bias-free reasoning collectively with discouraging biased reasoning.
-As shown in Figure 2 right panel, we first create two gender-agnostic sentences by replacing the gender pronoun with the two occupations that appeared in the original sentence.
-Then we ask the model a factual question: which sentence is more likely to happen in real life, as this question does not contain any gender-related information.
-We then prepend its answer (to the factual question) before asking the original question.
-This allows us to distill LLM’s non-gender-related world knowledge (Strategy III) and nudge it to explicitly reason with this fact during coreference resolution (Strategy I).
-On top of encouraging bias-free reasoning (Strategies I and III), we also prompt that both occupations are equally likely to be male or female to counteract existing selection bias (Strategy II), discouraging biased reasoning.
-One caveat of utilizing LLM’s world knowledge is that the performance of Reduce + Fact will increase as the capabilities of the LLM grow, since better world knowledge will further help LLM answer both factual and original questions.
+ICL (16-shot) 46.81 94.57 47.76 45.18 92.81 47.63 73.68 92.27 18.59 88.87 98.10 9.23
+DDP (ours) 73.27 73.95 0.68 72.73 84.67 11.94 74.08 75.17 1.09 94.57 96.74
+Type II    Default 85.01 97.97 12.96 94.28 98.98 4.70 93.77 97.46 3.68 97.59 99.49 1.91 COT (zero shot) 69.12 86.79 17.66 93.90 98.35 4.45 95.30 99.62 4.32 97.97 99.62 1.65 ICL (16-shot) 94.41 99.62 5.21 95.43 98.98 3.56 94.03 97.84 3.81 98.35 99.87 1.52 DDP (ours) 83.23 84.24 1.02 92.12 94.41 2.29 82.34 85.26 2.92 99.62 99.75 0.13
 4.1.2 Our Results (WinoBias)
-We compare Reduce + Fact with the three baseline approaches on both coreference tasks and summarize our results in Table 3.2.
-We expect all LLMs to achieve higher accuracies on sentences containing pro-stereotypical pronouns (pro sentences) compared to these with anti-stereotypical pronouns (anti sentences) due to historical biases in the training data.
+We compare DDP with the three baseline approaches on both coreference tasks and summarize our results in Table 4.1.1.
+Because of the existence of historical biases in LLMs’ training data, we expect all LLMs to score higher accuracies on sentences containing pro-stereotypical pronouns (pro sentences) compared to these with anti-stereotypical pronouns (anti sentences).
 Coreference Resolution Requiring World Knowledge
@@
 The method Zero-shot COT marginally reduces the bias, as seen in the smaller gaps, but its performance is lower on both pro and anti scenarios for GPT-3 and GPT-3.5 when compared with Default, and it only marginally improves the general performance when applied to more capable ones (Claude 2 and GPT-4).
-For ICL with counterfactual examples, although the gap becomes larger for less capable models (GPT-3 and GPT-3.5), it can further reduce biases with more capable LLMs (Claude 2 and GPT-4), especially for GPT-4 (with a gap of 9.23%).
-Remarkably, Reduce + Fact, which encourages bias-free reasoning while discouraging biased reasoning, substantially decreases the bias across all LLMs.
-The achieved minimal gaps, with GPT-4 exhibiting a mere 2.17% gap and 94.57% accuracy on anti sentences, demonstrate significantly greater effectiveness compared to other approaches.
-Coreference Resolution Only with Syntactic Information
-For Type II tasks, coreference can be resolved only with syntactic cues.
-All models generally show higher accuracies with smaller gaps.
-However, since there exists a shortcut in the Type II task – the correct answer is always the second entity in the sentence – a smaller gap may not necessarily indicate less bias.
-Still, Reduce + Fact outperforms the three baselines, particularly with GPT-4, achieving a near-negligible gap of 0.13%, suggesting that our method is highly effective in encouraging models to utilize non-gendered cues for coreference resolution.
+For ICL with contrastive examples, although the gap becomes larger for less capable models (GPT-3 and GPT-3.5), it can further reduce biases with more capable LLMs (Claude 2 and GPT-4), especially for GPT-4 (with a gap of 9.23%).
+Remarkably, Dual Directional Prompting, which employs multiple strategies to encourages fact-based reasoning substantially decreases the bias across all LLMs.
+They achieved minimal gaps, with GPT-4 exhibiting a mere 2.17% gap and 94.57% accuracy on anti sentences, demonstrate significantly greater effectiveness compared to other approaches.
+Coreference Resolution with Only Syntactic Information
+For Type II tasks, coreference can be resolved with only syntactic cues; thus, all models generally show higher accuracies with smaller gaps.
+However, since there exists a shortcut in the Type II task (the correct answer is always the second entity in the sentence), a smaller gap may not necessarily indicate less bias.
+Still, DDP outperforms the three baselines, particularly with GPT-4, achieving a near-negligible gap of 0.13%, suggesting that our method is highly effective in encouraging models to utilize gender-agnostic world knowledge for coreference resolution.
 Summary
-These experimental results suggest prompt designs that encourage bias-free reasoning and discourage biased reasoning are effective at mitigating gender biases in LLMs by directing them to rely more on non-gender-related world knowledge (Strategies I and III) and less on gender shortcuts (Strategy II), thus promoting fairer and less biased responses.
+These experimental results suggest prompt designs that direct LLMs to rely more on gender-agnostic world knowledge (Strategy I) and less on gender shortcuts (Strategy II), thus promoting fairer and less biased decisions.
 Also, the performance gap between pro sentences and anti sentences decreases as the LLMs become more capable, which may indicate that LLMs are less prone to assign occupations with stereotypical gender pronouns as their general (reasoning) capabilities grow.
-Figure 5: Performance comparison on Discrim-Eval across three demographics.
-The bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
-Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning).
-Encouraging bias-free reasoning universally decreases the relative gap when added with methods that reduce biased reasoning.
-4.1.3 Ablation studies
-We further conduct ablation studies to individually examine the effectiveness of the two types of approaches, i.e., encouraging bias-free reasoning and discouraging biased reasoning.
-We divide the prompt Reduce + Fact into two parts: the Reduce Only part prompting both men and women can perform either occupation equally, and the Fact Only part telling the model which sentence(s) it regards to be more viable in real life.
-Moreover, we categorize each LLM’s response into four groups: (i) TT: where the LLM answers both the factual question and the original question correctly, (ii) TF: where the LLM only answers the factual question correctly, (iii) FT: where the LLM only answers the original question correctly, and (iv) FF: where the LLM answers wrongly on both questions.
+4.1.3 Ablation studies (WinoBias)
+We further conduct ablation studies to individually examine the effectiveness of the two types of approaches, i.e., encouraging fact-based reasoning and discouraging biased reasoning.
+We divide the prompt Dual Directional Prompting into two parts: the Counteract Only part prompting that both occupations are equally likely to be male and female (Strategy II), and the Fact Only part informing the model which sentence(s) it regards to be more likely in real life (Strategy I).
+Moreover, we categorize each LLM’s response into four groups: (i) TT: where the LLM correctly answers both the base question and the original question, (ii) TF: where the LLM only answers the base question correctly, (iii) FT: where the LLM only answers the original question correctly, and (iv) FF: where the LLM answers wrongly on both questions.
 The above categorization allows us to attribute the model’s coreference mistakes into two categories: the mistake caused by its non-gender-related world knowledge (FF), and the mistakes caused by its gender bias (TF).
@@
 We divide the models’ responses into 4 categories to better understand the cause of their errors and success.
-$F$ denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge.
+TT denotes the (%) of examples where LLM answers both the base question and the original question correctly, FF denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge.
+TF denotes the coreference errors caused by gender bias as only base questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the base questions get wrong but original questions are correctly answered.
+More details about the settings can be found in Appendix C.1.
+GPT-4     DDP 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Fact Only 91.45 95.93 5.29 0.14 0.54 0.81 2.71 3.12 Counteract Only 70.69 88.20 25.78 8.28 0.95 1.76 2.58 1.76 Default 69.47 87.25 26.87 8.68 1.36 2.44 2.31 1.63
+As shown in Table 4.1.3, separating prompting strategies into Fact Only and Counteract Only, we observe that Fact Only still performs quite well, but Counteract Only leads to a significant decrease in performance.
+This suggests that while counteracting existing (selection) bias is important, providing explicit instructions that promote fact-based reasoning is crucial for mitigating bias in LLMs.
+Moreover, comparing the performance of GPT-4 with GPT-3.5, we can see that the GPT-4’s improvements mostly stem from better (non-gender-related) world knowledge (making way less mistakes on base questions) and self-consistency (less mistakes in the TF group).
+More ablation studies can be found in Appendix D.1, where we also design prompts to control the extent to which we counteract the existing bias in LLMs.
+4.2 Social Bias: BBQ
+The Bias Benchmark for QA (BBQ) (Parrish et al., 2021) is a dataset designed to evaluate bias for question-answering tasks.
+It highlight attested social biases against people belonging to protected classes along nine social dimensions relevant to U.S.
+English-speaking contexts.
+Besides common social dimensions such as age, gender, and race, BBQ also contains social categories such as physical appearance and socio-economic status where bias may be much more subtle.
+Each question in BBQ contains a context and 2 involved entities or persons, where the model needs to decide among 3 choices: one of the two entities, or unknown.
+4.2.1 Experimental Settings (BBQ)
+For our experiments, we consider the disambigous setting in BBQ where we test whether the model’s biases override a correct answer choice given an adequately informative context.
+In this setting, the answer is always one of the two entities rather than unknown.
+There are over 16,000 examples under this setting, and we measure the performance of GPT-4 across these nine social categories.
+Baseline Approaches
+We use similar baselines as in WinoBias (Section 4.1.1), where we compare DDP with Default and the stronger baseline ICL with contrastive examples.
+We change the number of ICL examples to 8 to match the settings in Si et al.
+(2022).
+Dual Directional Prompting on BBQ
+is also similar to what we defined on WinoBias.
+To encourage fact-based reasoning, we follow Figure 1(b) to first convert an original question into a base question where we remove social category information by using neutral references (i.e., Person X and Person Y) to label the two entities in the given context.
+Then we ask the LLM the base question and prepend its answer before asking the original question (Strategy I).
+To discourage biased reasoning, we ask it not to use the information related to the underlying social category when making the decision (Strategy III).
+4.2.2 Our Results (BBQ)
+In Table 3, we can see that on common social categories where biases may already be well-regulated (e.g., age, gender, and race), even the Default method has a reasonably good performance.
+Yet, on social categories where bias may be more subtle (e.g., physical appearance, religion, socio-economic status, and sexual orientation), the model has a much lower accuracy when no additional prompts are used.
+On these categories both ICL and DDP significantly improved the performance over Default, and DDP achieves the highest performance across 8 out of 9 social categories.
+Table 3: Performance comparison of various debiasing methods on BBQ. We measure how prompting-based debiasing methods help GPT-4 predict the correct choice given an adequately informative context.
+Accuracy (%) $\uparrow$ Age Disability Status Gender Identity Nationality Physical Appearance Race Ethnicity Religion SES Sexual Orientation
+Default 94.78 89.46 89.70 97.99 72.34 90.23 74.17 68.36 68.98
+ICL (8-shot) 99.95 97.17 98.24 99.68 82.11 97.09 93.00 87.15 91.20
+DDP (ours) 99.95 95.76 99.33 100.00 88.71 99.07 97.67 97.58 97.69
+5 Conclusion
+Our work presents a causality-guided, prompting-based debiasing framework for LLMs’ decisions.
+The framework tackles social biases and provides a novel perspective on how social category information can affect LLMs’ decisions via different causal pathways.
+Leveraging these causal insights, we propose principled, prompting-based debiasing strategies that utilize selection mechanisms to regulate the influence of bias.
+These strategies aim to discourage biased reasoning and encourage fact-based reasoning, unifying existing prompt-based debiasing methods and addressing their critical gap in promoting fact-based reasoning.
+Besides strong empirical results, we also prove that bias can be completely removed from LLMs’ decisions when the objectives in all three strategies are satisfied.
+For future work, our proposed Dual Directional Prompting (DDP) method may also be used to identify pairs of positive (unbiased) and negative (biased) responses to learn a reward model (utilizing the intuition that an unbiased response should align with the model’s base decision).
+Broader Impact
+In this paper, we present a causality-based LLM output debiasing framework.
+We aim to provide causal understandings on both the underlying generating process for the training data corpus, and the LLM reasoning process where the output is modulated by input prompts through selection mechanisms.
+Our proposed framework serves as a general solution to promote fairness in pretrained LLMs without accessing their model parameters.
+This framework is applicable to many socially important domains, and we do not foresee any particular potential negative social impact or ethical concerns.
+Acknowledgment
+We would also like to acknowledge the support from NSF Award No.
+2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Quris AI, Florin Court Capital, and MBZUAI-WIS Joint Program.
+ZT is supported by the National Institute of Justice (NIJ) Graduate Research Fellowship, Award No.
+15PNIJ-24-GG-01565-RESS.
+Appendix A Related Works
+In this section, we provide detailed discussions on related works.
+We consider the combinations of very related topics including causality, algorithmic fairness, and LLM reasoning.
+In Section A.1, we reviewed existing literature on how they address selection.
+In Section A.2, we consider causal notions of fairness that do not specifically pertain to the LLM context.
+In Section A.3, we consider existing efforts to draw connections between causality and LLM reasoning.
+In Section A.4, we consider previous works on LLM debiasing.
+In Section A.5, we consider previous works that involve all three topics.
+A.1 Selection mechanisms
+Previous literature has investigated selection mechanisms from diverse perspectives.
+For instance, the influence of selection bias on statistical inference in economic and sociological studies (Heckman, 1979; 1990; Winship & Mare, 1992), causal discovery when there are selection variables and latent common causes (Spirtes et al., 1995; Zhang, 2008), the identification and estimation of functional causal models when selection exists (Zhang et al., 2016), the identifiability of causal effect in the presence of selection bias from the graphical condition perspective (Bareinboim & Pearl, 2012; Bareinboim & Tian, 2015; Correa et al., 2019) and from the potential outcome perspective (Hernán & Robins, 2020), and the identification of the existence of selection bias from observational data under certain functional assumptions (Kaltenpoth & Vreeken, 2023).
+Recent work also demonstrates the identification of selection structures in sequential data (Zheng et al., 2024).
+A.2 Causality and Fairness
+It has been recognized in the algorithmic fairness literature that causality provides a unique tool to facilitate a better understanding of the data-generating process, and therefore, more effective bias quantifications and mitigations (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019; Nabi et al., 2022; von Kügelgen et al., 2022; Tang et al., 2023a).
+Previous causal fairness literature has considered notions based on estimating or bounding various kinds of causal effects (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019), and also the causal modeling of the dynamics as well as the long-term implications of bias mitigation strategies (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a; b).
+Because of the relatively abstract definition of the variable or node in the language context, previous approaches for characterizing and enforcing causal fairness are not directly applicable in LLM debiasing tasks.
+That being said, as we have demonstrated in our causality-guided debiasing framework, causal understandings of the involved data-generating processes help identify effective debiasing strategies that are both intuitively clear and theoretically grounded.
+A.3 Causality and LLMs
+The intersection between causality and LLMs has drawn increasing attention.
+Zhang et al.
+(2023) considers three types of causal questions and aims to evaluate LLMs’ abilities to identify causal relations, discover new knowledge from data, and quantitatively estimate the consequences of actions.
+Kıcıman et al.
+(2023) investigate LLMs’ abilities to perform causal reasoning and solve covariance-/logic- based causal questions.
+They also study the failure modes of LLMs and provide techniques to interpret the model robustness.
+Jin et al.
+(2023) propose a benchmark data set for evaluating LLMs’ causal inference capabilities via the task of determining causal relationships from a set of correlational statements.
+This line of research focuses on complex causal reasoning abilities in general settings, without specific attention to potential fairness violations.
+In comparison, our causality-guided debiasing framework does not involve assumptions/requirements on LLMs’ general-purpose causal reasoning capabilities.
+We adopt a rather mild assumption that a well-trained and well-aligned LLM captures the dependence pattern in the training data and that such a pattern is internalized and utilized during reasoning.
+A.4 Debiasing Language Models
+There is a large amount of work discussing bias and fairness in the context of language models (LMs) Bordia & Bowman (2019); Liang et al.
+(2021); Abid et al.
+(2021); Wang et al.
+(2023a); Liu et al.
+(2023); Ray (2023); Rozado (2023), and our investigation lies on debiasing techniques with causal understandings of the sources of biases.
+For debiasing approaches in the context of LMs, there are proposals involving direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), incorporating Reinforcement Learning with Human Feedback (RLHF) to better align the models with human values (Ouyang et al., 2022; Bai et al., 2022; Yao et al., 2023), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023; Furniturewala et al., 2024).
+We focus on prompting-based techniques and identify principles for prompt designs to steer LLMs toward unbiased responses by a).
+reducing biased reasoning and b).
+encouraging fact-based reasoning.
+We provide demonstrations of how we can employ the above two principles.
+Works on LLMs reasoning such as Wei et al.
+(2022); Zheng et al.
+(2023); Furniturewala et al.
+(2024) can also be incorporated into our framework to encourage fact-based reasoning(e.g., the Implication Prompting proposed in Furniturewala et al.
+(2024) can be viewed as an implicit approach encouraging bias-free reasoning).
+Though prompting-based debiasing methods become the only viable choice with black-box access, when we do have access to the model’s weights and enough compute resources to fine-tune the model, direct or instruction-based fine-tuning should be conducted first.
+Such approaches enable the integration of bias mitigation directly into the model parameters and can address the model’s intrinsic biases.
+A.5 Debiasing LLMs from Causal Perspectives
+Most closely related literature considers LLM debiasing strategies from causal perspectives.
+Vig et al.
+(2020) utilize causal mediation analysis and consider neuron-level intervention to investigate the instantiation of gender bias in Transformer-based language models (Vaswani et al., 2017).
+Zhang et al.
+(2024) also employs a similar view where they consider the chain-of-thought generated by LLMs as a mediator variable between prompt and answer.
+Wang et al.
+(2022) proposed a causal graph for relation extraction, and their counterfactual analysis lies in analyzing the change in the output relation distribution before and after removing the textual context.
+Zhou et al.
+(2023) propose Causal-Debias to mitigate the unwanted stereotypical association by fine-tuning pretrained language models.
+The causal model they consider involves four variables: label-relevant factor, bias-relevant factor, raw sentence, and ground-truth label.
+Wang et al.
+(2023b) pay special attention to entity bias and propose a specific structural causal model (SCM) for easier parameter estimations, such that the intervention-based mitigation strategy can be carried out.
+Their causal model involves four variables: entity, raw text, LLM input, and LLM output.
+In comparison, we provide detailed causal modelings at a sub-sentence level, considering both the training corpus generating process and the LLM reasoning process.
+Furthermore, our framework explicitly models the interplay between internal representations and external inputs through selection mechanisms, providing a clearer picture regarding possible strategies to debias LLM outputs more effectively.
+Our approach does not require white-box access or the ability to perform interventions, making our prompting-based framework applicable to a variety of practical scenarios.
+Appendix B Further Illustrations and Discussions of Our Framework
+In this section, we provide further illustrations on the debiasing strategies identified in Section 3.3.
+(c) LLM reasoning process (annotated with selection mechanisms specified by prompts)
+(d) LLM reasoning process (Strategy I: nudge towards social-agnositic fact)
+(e) LLM reasoning process (Strategy II: alternative fact counteracts existing selection bias)
+(f) LLM reasoning process (Strategy III: nudge aware from social-salient text)
+(g) LLM reasoning process (all strategies combined)
+Figure 4: Additional illustrations on debiasing strategies.
+In Figure 4, we use light coral (blue) to highlight unregulated (regulated) information flows from social category representations to LLM outputs in the internal reasoning process.
+We use mix-colored highlight to denote partially regulated information flow, e.g., the edge from social-salient text representation to LLM potential output in Figure 3(f).
+Comparing Figure 3(g) (all strategies combined) with Figures 3(d)3(f) (strategies applied individually), we can see that collectively, the strategies address the social bias in language more comprehensively.
+Here by “regulated,” we are referring to the constraint over the association between social category information representation and the potential output.
+We would like to note that our causal models can naturally handle situations where certain kinds of neutral dependence patterns involving demographic information are not necessarily considered problematic (Tang et al., 2024).
+Recent benchmark suite by Wang et al.
+(2025) also explicitly argues that group difference awareness matters when measuring desired group discrimination in LLMs (e.g., in legal contexts).
+We present relevant experimental results in Section D.1.
+We would also like to note that our causality-guided framework is not limited to situations where social category information is explicitly involved.
+One can adapt our framework to the specific practical setting by incorporating relevant nodes into causal graphs, thereby identifying the most suitable debiasing strategies therein.
+Appendix C Experimental Details
+Pretrained LLMs
+For the GPT models used in our experiments on WinoBias and Discrim-Eval, we consider snapshots from June 13th, 2023 where the knowledge cut-off time is Sep 2021.
+Since the legacy GPT-3 model (a.k.a., text-davinci-003) is no longer supported when we conduct the experiments, we use the model “gpt-3.5-turbo-instruct” instead as it has similar capabilities as GPT-3 era models.
+The Mistral-7B model we use in our experiments is the improved instruction fine-tuned version (a.k.a., “Mistral-7B-Instruct-v0.2”).
+For Mistral-7B, we choose the instruction-finetuned version over the base version as the effectiveness of prompting strategies in regulating biased pathways is strongly tied to the model’s ability to follow instructions.
+We are unable to use Claude 2 for our experiments on the Discrim-Eval dataset because we cannot access the probabilities of generated tokens.
+All LLMs responses are obtained with a temperature of 0.
+For our experiments on BBQ data set, we use the latest GPT-4 version (i.e., gpt-4-turbo).
+Iterative Prompting
+The responses of LLMs may not follow a given format even with specific instructions, which makes answer extraction challenging.
+Moreover, we observe that adding instructions to restrict the format of responses may lower the quality of the generated outputs.
+Therefore, we apply 2-round iterative prompting in our experiments where we let the models generate freely and then ask them to summarize their answers in one or two words.
+This way allows us to obtain high answer qualities with an easy extraction process at the same time.
+C.1 Gender Bias: WinoBias
+For experiments on the WinoBias dataset, we combined both the training and test data for evaluation as there is no need to separate them when using prompting-based debiasing techniques.
+We also noticed that the original dataset contains 400 grammar mistakes across 3000+ sentences in total.
+We corrected these grammar mistakes as they may hinder the co-reference resolution tasks.
+We have provided the cleaned version in the supplementary materials.
+Since there are 571 examples where two or more models make the same mistakes on the factual questions.
+After looking into these mistakes, we find that some of the examples are indeed ambiguous in nature (e.g., "The construction worker stared at the teacher because [who] was angry").
+Therefore, we collected human evaluations to understand the inherent ambiguity in the WinoBias dataset.
+For each of the 571 examples, we collect 5 human responses on the factual questions (i.e., which sentence is more likely to happen in real life), and we ask the human annotator to choose from {Sentence A, Sentence B, and Equally likely}.
+Based on the human responses, we identified 60 examples (55 from the Type I task and 5 from the Type II task) where three or more annotators disagree with the ground truth answers or think both sentences in the factual questions are equally likely to happen in real life.
+We removed these 60 examples during our evaluation, and we will include the human evaluation results in our codebase.
+Also, for instances where the model was prompted more than three times but still failed to provide a definitive choice (e.g., when presented with options A and B but refused to select either), those refusal examples were removed from our calculations (e.g., in Tables 4.1.3C.2).
+This approach ensures the reported percentages accurately reflect the cases where the model made a conclusive decision.
+The ICL baseline on the WinoBias dataset is from Si et al.
+(2022), where they constructed various sets of ICL examples (e.g., examples exclusively from anti-bias scenarios for Type I questions) and compared their effectiveness in debiasing LLMs’ outputs.
+Their findings reveal while using ICL examples from a specific category enhances performance for that category (e.g., pro-bias examples improve pro-bias predictions), a balanced set of pro-bias and anti-bias examples proved most effective in reducing bias.
+Therefore, this balanced strategy is one of the baselines we compare against.
+C.2 Demographic Bias: Discrim-Eval
+The Discrim-Eval dataset contains 70 diverse decision scenarios and $9\times 3\times 5\times 70=9450$ individual decision questions which includes all combinations of [AGE] $\in$ [20, 30, 40, 50, 60, 70, 80, 90, 100], [GENDER] $\in$ [male, female, non-binary] and [RACE] $\in$ [white, Black, Asian, Hispanic, Native American].
+To measure the corresponding bias in each demographic category, we reconstruct the dataset by extracting the base scenario which does not contain any demographic information (e.g., we replace all pronouns with the anaphoric reference to avoid leaking the gender information).
+We then ask the model to decide on each of the 70 base scenarios.
+There are (1/11/1/2) scenarios where (Mistral 7B/GPT-3/GPT-3.5/GPT-4) refuses to answer or does not output a Yes answer, and we removed these scenarios correspondingly when evaluating these LLMs.
+Table 4: Error Analysis on WinoBias Type II coreference task.
+We divide the models’ responses into 4 categories to better understand their success and failure cases.
+TT denotes the (%) of examples where LLM answers both the factual question and the original question correctly, FF denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge.
 TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
-GPT-4   Reduce + Fact 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Fact Only 91.45 95.93 5.29 0.14 0.54 0.81 2.71 3.12 Reduce Only 70.69 88.20 25.78 8.28 0.95 1.76 2.58 1.76 Default 69.47 87.25 26.87 8.68 1.36 2.44 2.31 1.63
-As shown in Table 4.1.3, when we separate the prompting strategies into Fact Only and Reduce Only, we observe that Fact Only still performs quite well, but Reduce Only leads to a significant decrease in performance.
-This suggests that while discouraging biased reasoning is important, providing explicit instructions that promote bias-free reasoning is crucial for mitigating bias in LLMs.
-Moreover, comparing the performance of GPT-4 with GPT-3.5, we can see that the GPT-4’s improvements mostly stem from better (non-gender-related) world knowledge (making way less mistakes on factual questions) and self-consistency (less mistakes in the TF group).
-More experimental results can be found in Appendix D.1, where we design prompts to control the degree to which we steer LLMs away from biased reasoning.
-4.2 Demographic Bias: Discrim-Eval
+GPT-4     DDP 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25 Fact Only 99.36 99.87 0.25 0.00 0.13 0.00 0.25 0.13 Counteract Only 98.98 99.49 0.76 0.13 0.00 0.38 0.25 0.00 Default 98.73 99.75 0.89 0.13 0.25 0.00 0.13 0.13
+Table 5: Error analysis of Dual Directional Prompting  on Type I and Type II questions in WinoBias.
+&82.47 82.34 1.27 1.40 0.76 1.91 15.50 14.36     GPT-3.5     Type I 70.15 78.70 10.58 2.04 2.58 5.97 16.55 13.16 Type II 88.69 89.71 1.14 0.13 3.43 4.70 6.73 5.46     Claude     Type I 73.00 73.00 2.99 2.99 1.09 2.17 22.93 21.85 Type II 79.80 81.19 6.61 5.21 2.54 4.07 11.05 9.53     GPT-4     Type I 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Type II 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25
+Table 6: Ablation study employing Strategy III to reduce biased reasoning with Claude 2 on WinoBias
+[…]
+DDP (Strategy III) 55.81 83.59 27.78
+DDP (Strategies II+III) 75.13 79.29 4.16
+DDP (Strategies I+II) 74.08 75.17 1.09
+DDP (Strategies I+II+III) 78.91 79.92 1.01
+Table 7: Adjusting the levels of counteracting existing selection bias on GPT-3.5 with Dual Directional Prompting.
+We investigated how the level of counteracting stereotype impacts the model performance by adjusting the Fact part in the input prompt.
+For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time.
+Appendix D Additional Experimental Results
+D.1 Gender Bias: WinoBias
+We include additional ablation studies on the WinoBias dataset, which include a detailed error analysis on WinoBias Type II coreference task (Table C.2), a detailed error analysis of Dual Directional Prompting with 4 LLMs on Type I and Type II tasks (Table C.2), an ablation study related to Strategy III—nudge aware from social-salient text—(Table 6), and an ablation study on adjusting the levels of biased reasoning (Table C.2) in the prompt design.
+In Section C.2, we show the detailed results on both Type I and Type II coreference tasks across 4 LLMs.
+As we can see, our method has bigger improvements on models with better world knowledge as.
+models with worse world knowledge could limit our method to reaching its full capacity.
+Ablation study on adjusting the levels of counteracting existing selection bias
+In Section C.2, we investigated how different levels of enforcing Strategy II (counteract existing selection bias) impacts the model debiasing performance by adjusting the Fact part in the input prompt.
+For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time; while a level of 50% indicates the following prompt: Assume that the physician can be male $50\%$ of the time and female $50\%$ of the time, and assume that the secretary can be male $50\%$ of the time and female $50\%$ of the time.
+We observe that as the level of anti-stereotype goes down, the errors caused by the use of the gender shortcut increase (TF increases).
+In addition, by soft adjustment of reducing biased reasoning, we provide not only flexible tuning strategies for the best model performance but also a chance to dive into the underlying reasons for the error.
+D.2 Demographic Bias: Discrim-Eval
 The Discrim-Eval data set (Tamkin et al., 2023) encompasses a collection of scenarios, each depicting a hypothetical case where a decision is required, such as approving a loan or issuing press credentials.
@@
 The individuals can be characterized by three demographic factors: age (within a range of 20 to 100, in 10-year intervals), gender (male, female, or non-binary), and race (White, Black, Asian, Hispanic, or Native American).
-4.2.1 Experimental Settings (Discrim-Eval)
+D.2.1 Experimental Settings (Discrim-Eval)
 The scenarios in the Discrim-Eval data set are designed such that an affirmative decision (Yes) is always favorable.
@@
 (2023).
-To amplify bias-free reasoning (Fact), we first remove all demographic information in each scenario and ask the LLM to make a Yes or No decision based on this base scenario.
-Then we explicitly include its answer to the base scenario in the prompt and ask the model to decide on the original scenario which contains specific demographic information (e.g., age 20).
-4.2.2 Our Results (Discrim-Eval)
+To amplify fact-based reasoning (Fact), we first remove all demographic information in each scenario and ask the LLM to make a Yes or No decision based on this base scenario (Strategy I).
+Then we explicitly include its answer to the base scenario in the prompt and ask the model to decide on the original scenario which contains specific demographic information.
+Figure 5: Performance comparison on Discrim-Eval across three demographics.
+The bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
+Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
+Encouraging fact-based reasoning universally decreases the relative gap when added with methods that reduce biased reasoning.
+D.2.2 Our Results (Discrim-Eval)
 Since the evaluation metric requires accessing the logit of each token, our experiments on Discrim-Eval are conducted on these four LLMs: Mistral 7B, GPT-3, GPT-3.5, and GPT-4.
 Figure 5 shows our experimental results on three demographic categories: gender, ethnicity, and age.
-Figure 6: GPT has an inherent bias towards aging. GPT-3’s default response (black) is more biased as age increases.
 Each bar in the figure represents the relative gap, which measures the disparity in the probability of making an affirmative decision (‘Yes’) for the most and least privileged groups within each demographic category, characterizing the relative extent of discrimination.
-The bars are color-coded according to the methods used, where lighter shades correspond to the methods that explicitly encourage bias-free reasoning (e.g.
+The bars are color-coded according to the methods used, where lighter shades correspond to the methods that explicitly encourage fact-based reasoning (e.g.
 purple bars denote Illegal and light purple bars denote Illegal + Fact).
@@
 While different prompts of discouraging biased reasoning show various degrees of effectiveness in reducing demographic bias, no single prompt stands out as the universally most effective option across all LLMs and demographic categories.
-However, encouraging bias-free reasoning (Fact) collectively with discouraging biased reasoning can further decrease the relative gap, compared to employing the approaches on their alone.
+However, encouraging fact-based reasoning (Fact) collectively with discouraging biased reasoning can further decrease the relative gap, compared to employing the approaches on their alone.
 This observation is consistent across 4 LLMs and 3 demographic categories.
@@
 Adverse Impact on Specific Demographic Groups
+Figure 6: GPT has an inherent bias towards aging. GPT-3’s default response (black) is more biased as age increases.
 We further examine the relative gaps for specific demographic features.
@@
 This suggests that GPT-3’s default responses may be more biased towards more senior individuals compared to junior ones.
-It is worth noting that the trend may potentially differ across LLMs, and we include additional plots for other LLMs on different demographic groups in Appendix D.2.
-5 Conclusion
-This paper presents a causality-guided and prompting-based LLM debiasing framework.
-In particular, we highlight the critical role of selection mechanisms in modeling data corpus bias and, more importantly, formulating how prompt designs can influence LLM outputs by specifying different selection conditions on its internal representations.
-Guided by causal understandings of such interplay, we identify principled prompting strategies for more effective debiasing.
-Our strong empirical results demonstrate the benefits of our framework, offering clear intuitions and theoretical foundations for effective debiasing approaches.
-Future work will naturally extend to acquiring bias-free knowledge and representations for LLMs.
-Broader Impact
-In this paper, we present a causality-based LLM output debiasing framework.
-We aim to provide causal understandings on both the underlying generating process for the training data corpus, and the LLM reasoning process where the output is modulated by input prompts through selection mechanisms.
-Our proposed framework serves as a general solution to promote fairness in pretrained LLMs without accesssing their model parameters.
-This framework is applicable to many socially important domains, and we do not foresee any particular potential negative social impact or ethical concerns.
-Acknowledgement
-This material is based upon work supported by the AI Research Institutes Program funded by the National Science Foundation (NSF) under AI Institute for Societal Decision Making (AI-SDM), Award No.
-2229881.
-This project is also partially supported by an Amazon Research Award (Fall 2022 CFP), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Apple Inc., KDDI Research Inc., Quris AI, and Infinite Brain Technology.
-Appendix A Related Works
-In this section, we provide detailed discussions on related works.
-We consider the combinations of very related topics including causality, algorithmic fairness, and LLM reasoning.
-In Section A.1, we consider causal notions of fairness that do not specifically pertain to the LLM context.
-In Section A.2, we consider existing efforts to draw connections between causality and LLM reasoning.
-In Section A.3, we consider previous works on LLM debiasing.
-In Section A.4, we consider previous works that involve all three topics.
-A.1 Causality and Fairness
-It has been recognized in the algorithmic fairness literature that causality provides a unique tool to facilitate a better understanding of the data-generating process, and therefore, more effective bias quantifications and mitigations (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019; Nabi et al., 2022; von Kügelgen et al., 2022; Tang et al., 2023a).
-Previous causal fairness literature has considered notions based on estimating or bounding various kinds of causal effects (Kilbertus et al., 2017; Kusner et al., 2017; Nabi & Shpitser, 2018; Nabi et al., 2019; Chiappa, 2019), and also the causal modeling of the dynamics as well as the long-term implications of bias mitigation strategies (Creager et al., 2020; Zhang et al., 2020; Tang et al., 2023a, b).
-Because of the relatively abstract definition of the variable or node in the language context, previous approaches for characterizing and enforcing causal fairness are not directly applicable in LLM debiasing tasks.
-That being said, as we have demonstrated in our causality-guided debiasing framework, causal understandings of the involved data-generating processes help identify effective debiasing strategies that are both intuitively clear and theoretically grounded.
-A.2 Causality and LLMs
-The intersection between causality and LLMs has drawn increasing attention.
-Zhang et al.
-(2023) considers three types of causal questions and aims to evaluate LLMs’ abilities to identify causal relations, discover new knowledge from data, and quantitatively estimate the consequences of actions.
-Kıcıman et al.
-(2023) investigate LLMs’ abilities to perform causal reasoning and solve covariance-/logic- based causal questions.
-They also study the failure modes of LLMs and provide techniques to interpret the model robustness.
-Jin et al.
-(2023) propose a benchmark data set for evaluating LLMs’ causal inference capabilities via the task of determining causal relationships from a set of correlational statements.
-This line of research focuses on complex causal reasoning abilities in general settings, without specific attention to potential fairness violations.
-In comparison, our causality-guided debiasing framework does not involve assumptions/requirements on LLMs’ general-purpose causal reasoning capabilities.
-We adopt a rather mild assumption that a well-trained and well-aligned LLM captures the dependence pattern in the training data and that such a pattern is internalized and utilized during reasoning.
-A.3 Debiasing Language Models
-There is a large amount of work discussing bias and fairness in the context of language models (LMs) (Bordia & Bowman, 2019; Liang et al., 2021; Abid et al., 2021; Wang et al., 2023a; Liu et al., 2023; Ray, 2023; Rozado, 2023), and our investigation lies on debiasing techniques with causal understandings of the sources of biases.
-For debiasing approaches in the context of LMs, there are proposals involving direct fine-tuning of model parameters (Kaneko & Bollegala, 2021; Garimella et al., 2021; Lauscher et al., 2021; Guo et al., 2022), modifying the decoding steps (Schick et al., 2021), incorporating Reinforcement Learning with Human Feedback (RLHF) to better align the models with human values (Ouyang et al., 2022; Bai et al., 2022; Yao et al., 2023), and prompting-based techniques (Si et al., 2022; Tamkin et al., 2023; Oba et al., 2023; Ganguli et al., 2023).
-We focus on prompting-based techniques and identify principles for prompt designs to steer LLMs toward unbiased responses by a).
-reducing biased reasoning and b).
-encouraging bias-free reasoning.
-We provide demonstrations of how we can employ the above two principles.
-Works on LLMs reasoning such as Wei et al.
-(2022); Zheng et al.
-(2023) can also be incorporated into with our framework to encourage bias-free reasoning.
-A.4 Debiasing LLMs from Causal Perspectives
-Most closely related literature considers LLM debiasing strategies from causal perspectives.
-Vig et al.
-(2020) utilize causal mediation analysis and consider neuron-level intervention to investigate the instantiation of gender bias in Transformer-based language models (Vaswani et al., 2017).
-Zhou et al.
-(2023) propose Causal-Debias to mitigate the unwanted stereotypical association by fine-tuning pretrained language models.
-The causal model they consider involves four variables: label-relevant factor, bias-relevant factor, raw sentence, and ground-truth label.
-Wang et al.
-(2023b) pay special attention to entity bias and propose a specific structural causal model (SCM) for easier parameter estimations, such that the intervention-based mitigation strategy can be carried out.
-Their causal model involves four variables: entity, raw text, LLM input, and LLM output.
-In comparison, we provide detailed causal modelings at a sub-sentence level, considering both the training corpus generating process and the LLM reasoning process.
-Furthermore, our framework explicitly models the interplay between internal representations and external inputs through selection mechanisms, providing a clearer picture regarding possible strategies to debias LLM outputs more effectively.
-Our approach does not require white-box access or the ability to perform interventions, making our prompting-based framework applicable to a variety of practical scenarios.
-(a) LLM reasoning process (annotated with selection mechanisms specified by prompts)
-(b) LLM reasoning process (Strategy I: nudge towards demographic-agnositic fact)
-(c) LLM reasoning process (Strategy II: counteract existing selection bias)
-(d) LLM reasoning process (Strategy III: nudge aware from demographic-aware text)
-(e) LLM reasoning process (all strategies combined)
-Figure 7: Additional illustrations on debiasing strategies.
-Appendix B Further Illustrations and Discussions of Our Framework
-In this section, we provide further illustrations on the debiasing strategies identified in Section 3.2.
-In Figure 7, we use light coral (blue) to highlight unregulated (regulated) information flows from demographic representations to LLM outputs in the internal reasoning process.
-We use mix-colored highlight to denote partially regulated information flow, e.g., the edge from demographic-aware text representation to LLM potential output in Figure 6(d).
-Comparing Figure 6(e) (all strategies combined) with Figures 6(b)6(d) (strategies applied individually), we can see that collectively, the strategies address the social bias in language more comprehensively.
-Here by “regulated”, we are referring to the constraint over the association between demographic information representation and the potential output.
-We would like to note that our causal models can naturally handle situations where certain kinds of neutral dependence patterns involving demographic information are not necessarily considered problematic (Tang et al., 2024).
-We present relevant experimental results in Section D.1.
-We would also like to note that our causality-guided framework is not limited to situations where demographic information is explicitly involved.
-One can adapt our framework to the specific practical setting by incorporating relevant nodes into causal graphs, thereby identifying the most suitable debiasing strategies therein.
-Appendix C Experimental Details
-Pretrained LLMs
-For the GPT models used in our experiments, we consider snapshots from June 13th, 2023 where the knowledge cut-off time is Sep 2021.
-Since the legacy GPT-3 model (a.k.a., text-davinci-003) is no longer supported when we conduct the experiments, we use the model “gpt-3.5-turbo-instruct” instead as it has similar capabilities as GPT-3 era models.
-The Mistral-7B model we use in our experiments is the improved instruction fine-tuned version (a.k.a., “Mistral-7B-Instruct-v0.2”).
-We are unable to use Claude 2 for our experiments on the Discrim-Eval dataset because we cannot access the probabilities of generated tokens.
-All LLMs responses are obtained with a temperature of 0.
-Iterative Prompting
-The responses of LLMs may not follow a given format even with specific instructions, which makes answer extraction challenging.
-Moreover, we observe that adding instructions to restrict the format of responses may lower the quality of the generated outputs.
-Therefore, we apply 2-round iterative prompting in our experiments where we let the models generate freely and then ask them to summarize their answers in one or two words.
-This way allows us to obtain high answer qualities with an easy extraction process at the same time.
-C.1 Gender Bias: WinoBias
-For experiments on the WinoBias dataset, we combined both the training and test data for evaluation as there is no need to separate them when using prompting-based debiasing techniques.
-We also noticed that the original dataset contains 400 grammar mistakes across 3000+ sentences in total.
-We corrected these grammar mistakes as they may hinder the co-reference resolution tasks.
-We have provided the cleaned version in the supplementary materials.
-Since there are 571 examples where two or more models make the same mistakes on the factual questions.
-After looking into these mistakes, we find that some of the examples are indeed ambiguous in nature (e.g., ”The construction worker stared at the teacher because [who] was angry”).
-Therefore, we collected human evaluations to understand the inherent ambiguity in the WinoBias dataset.
-For each of the 571 examples, we collect 5 human responses on the factual questions (i.e., which sentence is more likely to happen in real life), and we ask the human annotator to choose from {Sentence A, Sentence B, and Equally likely}.
-Based on the human responses, we identified 60 examples (55 from the Type I task and 5 from the Type II task) where three or more annotators disagree with the ground truth answers or think both sentences in the factual questions are equally likely to happen in real life.
-We removed these 60 examples during our evaluation, and we will include the human evaluation results in our codebase.
-C.2 Demographic Bias: Discrim-Eval
-The Discrim-Eval dataset contains 70 diverse decision scenarios and $9\times 3\times 5\times 70=9450$ individual decision questions which includes all combinations of [AGE] $\in$ [20, 30, 40, 50, 60, 70, 80, 90, 100], [GENDER] $\in$ [male, female, non-binary] and [RACE] $\in$ [white, Black, Asian, Hispanic, Native American].
-To measure the corresponding bias in each demographic category, we reconstruct the dataset by extracting the base scenario which does not contain any demographic information (e.g., we replace all pronouns with the anaphoric reference to avoid leaking the gender information).
-We then ask the model to decide on each of the 70 base scenarios.
-There are (1/11/1/2) scenarios where (Mistral 7B/GPT-3/GPT-3.5/GPT-4) refuses to answer or does not output a Yes answer, and we removed these scenarios correspondingly when evaluating these LLMs.
-Table 3: Error Analysis on WinoBias Type II coreference task.
-We divide the models’ responses into 4 categories to better understand their success and failure cases.
-$T$ denotes the (%) of examples where LLM answers both the factual question and the original question correctly, $F$ denotes the (%) of examples where both questions are answered incorrectly, indicating the coreference errors caused by the model’s world knowledge.
-TF denotes the coreference errors caused by gender bias as only factual questions are correctly answered, and FT indicates coreference success that may be due to gender shortcut since the factual questions get wrong but original questions are correctly answered.
-GPT-4   Reduce + Fact 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25 Fact Only 99.36 99.87 0.25 0.00 0.13 0.00 0.25 0.13 Reduce Only 98.98 99.49 0.76 0.13 0.00 0.38 0.25 0.00 Default 98.73 99.75 0.89 0.13 0.25 0.00 0.13 0.13
-Table 4: Error analysis of Reduce + Fact on Type I and Type II questions in WinoBias.
-82.47 82.34 1.27 1.40 0.76 1.91 15.50 14.36   GPT-3.5   Type I 70.15 78.70 10.58 2.04 2.58 5.97 16.55 13.16 Type II 88.69 89.71 1.14 0.13 3.43 4.70 6.73 5.46   Claude   Type I 73.00 73.00 2.99 2.99 1.09 2.17 22.93 21.85 Type II 79.80 81.19 6.61 5.21 2.54 4.07 11.05 9.53   GPT-4   Type I 94.44 96.07 1.76 0.27 0.14 0.68 3.66 2.99 Type II 99.62 99.62 0.13 0.00 0.00 0.13 0.25 0.25
-Table 5: Adjusting the levels of counteracting existing selection bias on GPT-3.5 with Reduce + Fact.
-We investigated how the level of counteracting stereotype impacts the model performance by adjusting the Fact part in the input prompt.
-For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time.
-Appendix D Additional Experimental Results
-D.1 Gender Bias: WinoBias
-We include additional ablation studies on the WinoBias dataset, which include a detailed error analysis on WinoBias Type II coreference task (Table C.2), a detailed error analysis of Reduce + Fact with 4 LLMs on Type I and Type II tasks (Table C.2), and an ablation study on adjusting the levels of biased reasoning (Table C.2) in the prompt design.
-In Section C.2, we show the detailed results on both Type I and Type II coreference tasks across 4 LLMs.
-As we can see, our method has bigger improvements on models with better world knowledge as.
-models with worse world knowledge could limit our method to reaching its full capacity.
-Ablation study on adjusting the levels of counteracting existing selection bias
-In Section C.2, we investigated how different levels of enforcing Strategy II (counteract existing selection bias) impacts the model debiasing performance by adjusting the Fact part in the input prompt.
-For example, a counteract level of 100% indicates the following prompt: Assume that the physician can be male $0\%$ of the time and female $100\%$ of the time, and assume that the secretary can be male $0\%$ of the time and female $100\%$ of the time; while a level of 50% indicates the following prompt: Assume that the physician can be male $50\%$ of the time and female $50\%$ of the time, and assume that the secretary can be male $50\%$ of the time and female $50\%$ of the time.
-We observe that as the level of anti-stereotype goes down, the errors caused by the use of the gender shortcut increase (TF increases).
-In addition, by soft adjustment of reducing biased reasoning, we provide not only flexible tuning strategies for the best model performance but also a chance to dive into the underlying reasons for the error.
-D.2 Demographic Bias: Discrim-Eval
-In Figure 8, Figure 9, Figure 10, Figure 11, we show in details the performance comparison on Discrim-Eval across three demographic categories across four LLMs (Mistral 7B, GPT 3, GPT 3.5, and GPT 4).
+It is worth noting that the trend may potentially differ across LLMs, and see additional plots for other LLMs on different demographic groups in Appendix D.2.3.
+D.2.3 Additional Plots
+In Figures 710, we show in detail the performance comparison on Discrim-Eval across three demographic categories across four LLMs (Mistral 7B, GPT-3, GPT-3.5, and GPT-4).
+The baseline setting is colored black.
+Across all four LLMs, we see that encouraging fact-based reasoning universally reduces the relative gap when used together with methods that discourage biased reasoning.
+Figure 7: Performance comparison on Discrim-Eval across three demographic categories with Mistral (7B).
 The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
-Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning).
-The baseline setting is colored black.
-Figure 8: Performance comparison on Discrim-Eval across three demographic categories in details on Mistral (7B).
+Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
+Figure 8: Performance comparison on Discrim-Eval across three demographic categories with GPT-3.
 The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
-Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning).
-Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
-Figure 9: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-3.
+Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
+Figure 9: Performance comparison on Discrim-Eval across three demographic categories with GPT-3.5.
 The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
-Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning).
-Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
-Figure 10: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-3.5.
+Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
+Figure 10: Performance comparison on Discrim-Eval across three demographic categories with GPT-4.
 The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
-Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning).
-Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
-Figure 11: Performance comparison on Discrim-Eval across three demographic categories in details on GPT-4.
-The height of the bar denotes the degree of discrimination by comparing the least privileged group with the most privileged group in a given demographic category (the higher the bar, the deeper the discrimination).
-Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify bias-free reasoning).
-Amplifying bias-free reasoning universally reduces the relative gap when added with methods that reduce biased reasoning.
+Different methods (prompt designs) are colored differently (lighter colors denote the ones that amplify fact-based reasoning).
+[…]
+[supp]